### v2

# Recursive Quantum Relaxation for Combinatorial Optimization Problems

Abstract
Quantum optimization methods use a continuous degree-of-freedom of quantum states to heuristically solve combinatorial problems, such as the MAX-CUT problem, which can be attributed to various NP-hard combinatorial problems. This paper shows that some existing quantum optimization methods can be unified into a solver that finds the binary solution that is most likely measured from the optimal quantum state. Combining this finding with the concept of quantum random access codes (QRACs) for encoding bits into quantum states on fewer qubits, we propose an efficient recursive quantum relaxation method called recursive quantum random access optimization (RQRAO) for MAX-CUT. Experiments on standard benchmark graphs with several hundred nodes in the MAX-CUT problem, conducted in a fully classical manner using a tensor network technique, show that RQRAO outperforms the Goemans–Williamson method and is comparable to state-of-the-art classical solvers. The codes will be made available soon.
1 Introduction
Figure 1: Concept of the proposed method, recursive quantum random access optimization (RQRAO). (a) One/two/three classical bits are embedded into one qubit using a quantum random access code (QRAC) [1, 2]. A total of two/four/eight variants of the information of the one/two/three classical bits are represented as a vector pointing to the corner of the line/square/cube inscribed in the Bloch sphere. These specific quantum states are called (one-qubit) magic states. (b) Using a high compression QRAC, the size of the Hilbert space corresponding to the given problem is reduced and each (multi-qubit) magic state is no longer necessarily orthogonal to the others. Note that only four of the $2^{6}$ magic states are illustrated. After optimizing the objective function defined on the reduced Hilbert space, the bit string is decoded by searching for the nearest magic state in the fidelity, which is called rounding. To accomplish the rounding without strong approximation, a recursive strategy similar to the RQAOA [3] is adopted.
The MAX-CUT problem is one of Karp’s 21 NP-complete problems [4] and is frequently used as a benchmark of combinatorial optimization algorithms because of its simplicity. Namely, given a graph that consists of nodes connected by weighted edges, the optimization variant of the MAX-CUT problem asks for the labeling of nodes into two classes such that the sum of the weighted edges connecting nodes with different labels is maximized. Obtaining good solutions for the MAX-CUT problem is in great demand, because various NP-hard combinatorial problems, e.g., partitioning, covering, and coloring problems, can be relaxed to the MAX-CUT problem by treating the constraints as penalty terms [5, 6]. Although the MAX-CUT problem is a discrete optimization problem, it can be translated to a continuous variant by treating the discrete solution in a probabilistic manner. Classically, such a continuous variant of the MAX-CUT problem is relaxed to a semidefinite program by approximating the probability [7]. However, the quantum optimization does not take such an approximation because the probabilistic characteristics are naturally incorporated into the quantum states. In the quantum MAX-CUT , all solutions are encoded in the eigenstates of the problem Hamiltonian, and then the quantum state is optimized such that the expected value of the problem Hamiltonian is maximized. From a theoretical viewpoint, it has been proved that the eigenstate with the maximum eigenvalue can be obtained by the infinite-level quantum approximate optimization algorithm (QAOA) [8, 9] with the appropriate parameters. However, in a noisy quantum device, it is intractable to carry out even a sufficiently large level QAOA, let alone an infinite level QAOA. For this reason, the field of combinatorial problems with quantum computing is dominated by the development of heuristics, that mainly focus on the development of the ansatz variants and improving the trainability of QAOA [10].
Rather than developing the QAOA ansatz, an alternative approach is to develop a non-diagonal Hamiltonian instead of the diagonal Ising-type one used in the QAOA. In [11], the MAX-CUT Hamiltonian is constructed using quantum random access codes (QRACs) [1, 2] (see Sec. 5.3 for a review). Using QRACs, the $n$-qubit quantum state can be used to encode $m$ binary variables for $m>n$. The QRAC Hamiltonian is still a 2-local Hamiltonian, and therefore it is QMA-hard to find the ground state [12], but the QRAC formulation has the advantage that the number of qubits can be reduced, making it easier to optimize the variational state.
Recently, [3] proposed the recursive QAOA (RQAOA), showing that the recursive application of level-1 QAOA gives an empirically higher approximation ratio than the naïve application of level-1 QAOA and the Goemans–Williamson (GW) algorithm [7], where the latter is the best generic classical algorithm with a theoretical guarantee for the MAX-CUT problem under the unique games conjecture [13]. This result is interesting because in [3], it was proved that the approximation ratio of the naïve level-1 QAOA cannot outperform the GW for some specific graphs. The application of the RQAOA has been extended to the MAX-$k$-CUT problem in [14], which also shows that the RQAOA outperforms the best-known classical approximation algorithms.
Here, the following questions naturally arise: Is it possible to apply the QRAC in the RQAOA framework? Furthermore, if it is possible, are there any benefits of using the QRAC in the RQAOA framework, such as improved approximation ratios? In this paper, we propose a general framework for the MAX-CUT problem that incorporates the QRAC [11] into the RQAOA [3] framework, namely recursive quantum random access optimization (RQRAO, see Fig. 1). Our contributions are summarized as follows:
  1. 1.
    We theoretically show that incorporating QRACs into the RQAOA framework is a natural expansion of [11] and [3], i.e., they are all uniformly expressed as a method that (i) maximizes the expectation of the problem Hamiltonian that all candidate solutions are embedded in using QRACs, and (ii) searches for the nearest embedding in the fidelity (Theorem 1).
  2. 2.
    We give numerical experiments to show the benefit of incorporating QRACs into the RQAOA framework. Demonstrations on an open graph dataset, Gset [15], were carried out using the tensor network simulation conducted on a classical computer. The results showed that the approximation ratio of the proposed method outperforms not only GW but also the methods from which inspiration was taken [11, 3], and is comparable to existing state-of-the-art classical heuristics [16].
The remainder of the paper is organized as follows: Section 2 presents the mathematical definition of the MAX-CUT problem and describes the RQAOA and QRAC in brief. Section 3 describes the proposed algorithm and its theoretical background. Section 4 presents the results of numerical experiments. Section 5 summarizes the related works. Section 6 concludes the paper.
2 Preliminaries
2.1 MAX-CUT Problem
As stated in Sec. 1, the goal of the MAX-CUT problem is to find binary node labels such that the sum of the weights of edges with different node labels is maximized. It is mathematically formulated as follows:
Definition 1 (MAX-CUT problem)
Given a graph $G=(V,E)$, where $V$ is a node set and $E$ is an edge set with edge weights $w_{jk}\in\mathbb{R}$ for $(j,k)\in E$, the MAX-CUT problem is defined as
$\max_{{\bf b}\in\{0,1\}^{|V|}}\texttt{CW}({\bf b})$ (1)
where
$\texttt{CW}({\bf b}):=\sum_{(j,k)\in E}w_{jk}\ \frac{1-(-1)^{b_{j}+b_{k}}}{2}$ (2)
is the cut weight corresponding to ${\bf b}$.
2.2 Recursive QAOA
Recursive QAOA (RQAOA) [3] is a variant of QAOA, which is an iterative method that determines the parity of one edge per iteration, then deletes a node and modifies a graph. Here, edge parity is defined as follows:
Definition 2 (Edge parity)
Given edge $(j,k)$, positive and negative edges are defined as $b_{j}=b_{k}$ and $b_{j}\neq b_{k}$, respectively, where $b_{i}\in\{0,1\}$ is an attribute of node $i$. Define edge parity as being a positive or negative edge.
In the first step of RQAOA, a candidate edge whose parity is to be determined is calculated using optimization. Let $|\psi({\boldsymbol{\theta}})\rangle$ be a QAOA ansatz with learnable parameters $\boldsymbol{\theta}$ and
$H_{\rm Ising}:=\sum_{(j,k)\in E}w_{jk}\,\frac{I-Z_{j}Z_{k}}{2}\$ (3)
be the Ising-type MAX-CUT Hamiltonian, where $Z_{j}$ is a Pauli $Z$ matrix corresponding to the $j$th qubit. The optimized parameters ${\boldsymbol{\theta}}^{*}$ are obtained by maximizing $\langle\psi({\boldsymbol{\theta}})|H_{\rm Ising}|\psi({\boldsymbol{\theta}})\rangle$. After optimization, the energy of edge $(j,k)\in E$ is calculated as
$\mathcal{E}_{jk}:=\langle\psi({\boldsymbol{\theta}}^{*})|Z_{j}Z_{k}|\psi({\boldsymbol{\theta}}^{*})\rangle.$ (4)
The candidate edge $(j^{*},k^{*})$ is defined as a maximal energy edge as
$(j^{*},k^{*})=\mathop{\rm argmax}_{(j,k)\in E}|\mathcal{E}_{jk}|.$ (5)
Then, the parity of edge $(j^{*},k^{*})$ is determined as
$\left\{\begin{array}[]{ll}b_{j^{*}}=b_{k^{*}}&\mathcal{E}_{j^{*}k^{*}}>0\\[14.22636pt] b_{j^{*}}\neq b_{k^{*}}&\mathcal{E}_{j^{*}k^{*}}<0\end{array}\right..$ (8)
Note that the probability that all edge energy is equal to zero ($\mathcal{E}_{j^{*}k^{*}}=0$) is negligibly small and can be ignored.
In the next step, the problem Hamiltonian is modified to satisfy Eq. (8) as a constraint. This modification can be interpreted as modifying the graph such that the number of its nodes is reduced by one, as shown in Fig. 2.
Figure 2: Graph modification procedure of RQAOA. (a) Initial graph. (b) The red edge indicates ${(j^{*},k^{*})=\mathop{\rm argmax}_{(j,k)}}|\mathcal{E}_{jk}|$, where $|\mathcal{E}_{jk}|:=|\langle\psi|Z_{j}Z_{k}|\psi\rangle|$. One of the nodes in edge $(j^{*},k^{*})$ is kept (the green node) while the other one is deleted (the red node). (c) After node removal. (d) Edges are re-connected to the retained node by multiplying ${\rm sgn}(\mathcal{E}_{j^{*}k^{*}})$ by its weight.
Node $j^{*}$ is deleted from the graph and edge $(j^{*},l)$ $(l\neq k^{*})$ is re-connected as $(k^{*},l)$. At the same time, edge weight $w_{k^{*}l}$ is updated to $w_{k^{*}l}+{\rm sgn}(\mathcal{E}_{j^{*}k^{*}})\cdot w_{j^{*}l}$ $(l\neq k^{*})$. As a result, one node is deleted from the graph by a single iteration. Subsequently, the variational state is re-initialized and optimized using the Hamiltonian defined with respect to the modified graph.
The number of nodes is gradually decreased by repeating the edge parity determination and graph size reduction. When the number of nodes becomes smaller than a threshold value $M$, the parities of the remaining nodes are determined by brute-force search.
2.3 Quantum Random Access Optimization
In the Quantum Random Access Optimization (QRAO) [11], binary node attributes are encoded in fewer qubits than the number of nodes using quantum random cccess code (QRAC) [1, 2]. For each node $j\in V$, binary node attribute $b_{j}\in\{0,1\}$ is randomly mapped to $(q^{(j)},P^{(j)})$, where $q^{(j)}$ is a qubit index and $P^{(j)}$ is a Pauli matrix with the following constraints:
$\left\{\begin{array}[]{ll}\displaystyle{P^{(j)}\neq P^{(k)}}&{\rm for\ \ \ }q^{(j)}=q^{(k)}\\[14.22636pt] \displaystyle{q^{(j)}\neq q^{(k)}}&{\rm\forall\ \ \ }(j,k)\in E\end{array}\right..$ (9)
In the quantum random access coding, there are variations that depend on the number of variants of the Pauli matrices. If only $Z$ is available, $P^{(j)}=Z$, it reduces to the Ising-type formulation, which we refer to $(1,1)$-QRAC. For $(2,1)$-QRAC, two variants of 1-local Pauli matrices, e.g., $P^{(j)}\in\{X,Z\}$, are available. For $(3,1)$-QRAC, $P^{(j)}\in\{X,Y,Z\}$ are available. Then, the relaxed MAX-CUT Hamiltonian, which we refer to as the $(m,1)$-QRAC Hamiltonian $(m=1,2,3)$, is defined as
$H_{m}:=\sum_{(j,k)\in E}w_{jk}\,\frac{I-mP_{\langle j\rangle}P_{\langle k\rangle}}{2},$ (10)
where $P_{\langle j\rangle}:=P^{(j)}_{q^{(j)}}$ is a 1-local Pauli matrix corresponding to the $q^{(j)}$th qubit. If the given graph is sparse, constraint Eq. (9)${}_{2}$ rarely restricts the assignment, and hence, the total number of qubits can be reduced by up to $1/m$th the number of nodes.
Let $\mu^{[q]}_{m}\in\mathbb{C}^{2\times 2}$ be the $q$th single qubit pure state in the form of a density matrix defined as
$\mu^{[q]}_{m}(\{b^{[q]}_{\bullet}\}):=\frac{1}{2}\Bigg{(}I+\frac{1}{\sqrt{m}}\sum_{P\in\mathfrak{P}}(-1)^{b^{[q]}_{P}}P\Bigg{)},$ (11)
where $\mathfrak{P}\subset\{X,Y,Z\}$, $|\mathfrak{P}|=m$, and $b^{[q]}_{P}\in\{0,1\}$ is a binary node attribute mapped to $(q,P)$.
Figure 3: Representation of the $(3,1)$-magic state in the Bloch sphere. The coordinates of the vertex of the cube inscribed in the sphere correspond to the embedded classical bits $(b^{[q]}_{X},b^{[q]}_{Y},b^{[q]}_{Z})$.
Because $\mu^{[q]}_{m}(\{b^{[q]}_{\bullet}\})$ is a one-qubit pure state, it can be drawn in the Bloch sphere (Fig. 3). For example, in the case of $m=3$, the states $\mu^{[q]}_{3}(\{b^{[q]}_{\bullet}\})$ for $(b^{[q]}_{X},b^{[q]}_{Y},b^{[q]}_{Z})\in\{0,1\}^{3}$ correspond to the vectors pointing the vertices of the cube inscribed in the Bloch sphere. The $n$-qubit product state of $\mu^{[q]}_{m}$ is defined as
$\mu_{m}({\bf b}):=\mu_{m}^{[n-1]}(\{b^{[n-1]}_{\bullet}\})\otimes\cdots\otimes\mu_{m}^{[0]}(\{b^{[0]}_{\bullet}\}).$ (12)
According to [11, Proposition 1], the following property holds.
${\rm tr}(H_{m}\,\mu_{m}({\bf b}))=\texttt{CW}({\bf b}).$ (13)
Of course, the $n$-qubit quantum state is not required to be a product state such as Eq. (12). If $\rho$ is assumed to be an arbitrary $n$-qubit state that includes the aforementioned product state, ${\rm tr}(H_{m}\,\rho)$ could be higher than ${\rm tr}(H_{m}\,\mu({\bf b}))$ when $m>1$, where $\rho$ may not be directly related to an optimal binary solution. This is why $H_{m}$ $(m>1)$ is called the relaxed Hamiltonian. Note that it has not been clarified what exactly ${\rm tr}(H_{m}\,\rho)$ represents when $\rho$ is an arbitrary state in [11], but this will be clarified in Corollary 1 in Sec. 3.2.
After maximizing ${\rm tr}(H_{m}\rho)$ with respect to $\rho$, the classical bit string ${\bf b}$ is decoded from the optimized $\rho$. This process is called rounding. In [11], two kinds of rounding were proposed, magic state rounding and Pauli rounding. Note that magic state rounding is a measurement method rather than a decoding method, and hence magic state rounding is henceforth referred to as random magic measurements. The random magic measurement is a projective measurement whose basis is the magic state selected uniformly at random. The detailed protocol is defined in Definition 3 of Supplementary Information A. Note that $(1,1)$-random magic measurement with $\mathfrak{P}=\{Z\}$ is equivalent to measurements on a computational basis. Pauli rounding is a decoding method based on the value of ${\rm tr}(P_{\langle j\rangle}\rho)$. The value of ${\bf b}$ is determined as $b_{j}=0$ if ${\rm tr}(P_{\langle j\rangle}\rho)>0$ and $b_{j}=1$ if ${\rm tr}(P_{\langle j\rangle}\rho)<0$. If ${\rm tr}(P_{\langle j\rangle}\rho)=0$, $b_{j}$ is determined as 00 or $1$ uniformly at random. In Sec. 3.2, we provide theorems to show the relationship between random magic measurements and Pauli rounding and reveal what assumption is made in Pauli rounding.
3 Proposed Method
The proposed method, named RQRAO, is introduced in this section.
Section 3.1 describes the concrete algorithm of the proposed method. In brief, the proposed method incorporates the $(m,1)$-QRAC Hamiltonian (Eq. (10)) into the RQAOA framework (Sec. 2.2). The edge parity is recursively determined based on the value of the edge energy, which is defined using the $(m,1)$-QRAC Hamiltonian. Two heuristics are also incorporated into the proposed method. One is the ensemble method of defining the edge energy, and the other determines several edge parities at once using the maximum spanning tree. The former improves the resulting cut weight whereas the latter reduces the runtime of the proposed method.
In Sec. 3.2, we show that not only the proposed method but also RQAOA and QRAO
  1. 1.
    try to maximize the expected cut weight using $(m,1)$-random magic measurements with respect to the quantum state, and
  2. 2.
    try to decode the bit string embedded in the nearest magic state in the view of the fidelity against the optimized quantum state.
3.1 Algorithm
The pseudocode of RQRAO is presented in Algorithm 1. The algorithm describes how the ensemble edge energy $\mathfrak{E}_{jk}$ of edge $(j,k)$ is calculated and how the edge parity is calculated from $\mathfrak{E}_{jk}$. The details are described below.
input : Graph $G=(V,E)$, number of ensembles $N$, scale factor $S$, brute-force search threshold $M$
output : MAX-CUT solution ${\bf b}^{*}$
initialize : $\mathcal{P}\leftarrow\emptyset$, $G^{\prime}=(V^{\prime},E^{\prime})\leftarrow G=(V,E)$
1 $w^{\prime}_{jk}\leftarrow w_{jk}+\xi,\ \ \ \xi\sim{\rm Uniform}([-10^{-5},10^{-5}])$;
2 while $|V^{\prime}|>M$ do
       /* Compute the ensemble edge energy */
3       for $t=1,\ldots,N$ do
4             Assign Pauli $P^{(t)}$ to each node $j\in V^{\prime}$ and make the $(m,1)$-QRAC Hamiltonian $H_{m}^{(t)}(G^{\prime})$;
5             $\theta^{(t)*}\leftarrow$ Result of maximization of $\langle\psi(\theta)|H_{m}^{(t)}(G^{\prime})|\psi(\theta)\rangle$;
6             ${\mathcal{E}^{(t)}_{jk}\leftarrow\langle\psi(\theta^{(t)*})|P^{(t)}_{\langle j\rangle}P^{(t)}_{\langle k\rangle}~{}|\psi(\theta^{(t)*})\rangle}$ for all $(j,k)\in E^{\prime}$;
7
8       end for
9      for $(j,k)\in E^{\prime}$ do
10             $\mu_{jk}\leftarrow\texttt{Mean}(\{\mathcal{E}_{jk}^{(t)}\})$;
11             $\sigma_{jk}\leftarrow\texttt{StandardDeviation}(\{\mathcal{E}_{jk}^{(t)}\})$;
12             ${\mathfrak{E}_{jk}\leftarrow}\mu_{jk}-{\rm sign}(\mu_{jk})\cdot{\rm min}(S\sigma_{jk},|\mu_{jk}|)$;
13
14       end for
      /* Compute the edge parity and modify the graph */
15       $E_{\rm S}\leftarrow\{(j,k,|\mathfrak{E}_{jk}|):(j,k)\in E^{\prime},|\mathfrak{E}_{jk}|>0\}$;
16       $E_{\rm T}\leftarrow\texttt{MaximumSpanningTree}(E_{\rm S})$;
17       for $(j,k)\in E_{\rm T}$ do
18             $\mathcal{P}\leftarrow\mathcal{P}\cup\{(j,k,{\rm sign}(\mathfrak{E}_{jk}))\}$;
19             $G^{\prime}=(V^{\prime},E^{\prime})\leftarrow\texttt{GetReducedGraph}(G^{\prime},(j,k))$
20       end for
21
22 end while
[…]
Algorithm 1 Algorithm for recursive quantum random access optimization (RQRAO).
Line 1: Add Noise to Edges
Add small perturbation noise to all edge weights to mitigate the isolated nodes problem (see Supplementary Information B).
Line 4: Define the Relaxed Hamiltonian
Assign 1-local Pauli $P^{(t)}$ to each node included in the current graph $G^{\prime}=(V^{\prime},E^{\prime})$, where superscript $(t)$ describes the $t$th trial. Note that $t$ is the sample index of the ensemble, not to be confused with the recursive iteration index.
The assignment is randomly carried out so as not to violate constraints Eq. (9). Then, the $t$th relaxed Hamiltonian is defined as
$H_{m}^{(t)}:=\sum_{(j,k)\in E^{\prime}}\frac{I-mP^{(t)}_{\langle j\rangle}P^{(t)}_{\langle k\rangle}}{2}w^{\prime}_{jk}.$ (14)
Line 5: Optimization
Define ansatz $|\psi({\boldsymbol{\theta}})\rangle$, where ${\boldsymbol{\theta}}\in\mathbb{R}^{K}$ are $K$ learnable parameters. The objective function is defined as
$\mathcal{L}^{(t)}({\boldsymbol{\theta}}):=\langle{\psi}(\boldsymbol{\theta})|H_{m}^{(t)}|{\psi}(\boldsymbol{\theta})\rangle.$ (15)
The optimized quantum state $|\psi({\boldsymbol{\theta}}^{*})\rangle$ is obtained by maximizing $\mathcal{L}^{(t)}({\boldsymbol{\theta}})$. Any optimization method can be used here.
Line 6: Compute the Edge Energy
The $t$th edge energy is computed as
$\mathcal{E}^{(t)}_{jk}:=\langle\psi({\boldsymbol{\theta}}^{*})|P^{(t)}_{\langle j\rangle}P^{(t)}_{\langle k\rangle}|\psi({\boldsymbol{\theta}}^{*})\rangle.$ (16)
The resulting edge energy strongly depends on both the assignment of 1-local Pauli $P^{(t)}$ and the reached quantum state, which is in general not a eigenstate corresponding to the maximum eigenvalue of $H_{m}^{(t)}$.
Lines 7–11: Compute the Ensemble Edge Energy
Using $N$ edge energies, $\{\mathcal{E}^{(t)}_{jk}\}^{N}_{t=1}$, the ensemble edge energy is calculated as
$\mathfrak{E}_{jk}:=\mu_{jk}-{\rm sign}(\mu_{jk})\cdot{\rm min}(S\sigma_{jk},|\mu_{jk}|)$ (17)
where $\mu_{jk}$ and $\sigma_{jk}$ are the mean and standard deviation of $\{\mathcal{E}^{(t)}_{jk}\}^{N}_{t=1}$, which are defined as $\mu_{jk}:=\frac{1}{N}\sum^{N}_{t=1}\mathcal{E}^{(t)}_{jk}$ and $\sigma_{jk}:=\sqrt{\frac{1}{N}\sum^{N}_{t=1}(\mathcal{E}^{(t)}_{jk}-\mu_{jk})^{2}}$, respectively. Equation (17) means that $\mathfrak{E}_{jk}$ has a nonzero value if $\mu_{jk}\pm S\sigma_{jk}$ have the same sign as $\mu_{jk}$.
Lines 12–16: Compute the Maximum Spanning Tree
After computing the ensemble edge energy $\mathfrak{E}_{jk}$, the parities of edges with nonzero $\mathfrak{E}_{jk}$ are determined according to the sign of $\mathfrak{E}_{jk}$. However, sometimes inconsistency arises. For example, this occurs when three negative edges form a cycle $\{(b_{0},b_{1}),(b_{1},b_{2}),(b_{2},b_{0})\}$, one of which is inconsistent because $(b_{0}\neq b_{1})\land(b_{1}\neq b_{2})\land(b_{2}\neq b_{0})$ cannot hold. To avoid this inconsistency, we adopt the maximum spanning tree to remove any cycles from the candidates.
Edges with nonzero $\mathfrak{E}_{jk}$ are collected and $|\mathfrak{E}_{jk}|$ are treated as edge weights. For each subgraph $E_{\rm S}$ made by this collection that is disconnected from the others, the maximum spanning tree $E_{\rm T}$ can be calculated in linear time in the number of edges by Prim’s algorithm [17] or by the randomized KKT algorithm [18]. The obtained tree $E_{\rm T}$ can be easily converted into a rooted tree by setting an arbitrary node as the root.
Lines 17–21: Modify the Graph
Now, we have candidates $E_{\rm T}$ that can be used to determine the parity and can be deleted from graph $G^{\prime}$ without inconsistency. The parity for each edge in $E_{\rm T}$ is determined from the leaf to the root one by one according to its sign. At the same time, the node on the leaf side is deleted and the edge weights are updated. The graph modification procedure is the same as the RQAOA, as depicted in Fig. 2.
Line 23: Brute-force Search
The recursive procedure is stopped when the number of nodes of $G^{\prime}$ is less than or equal to $M$. Here, the number of nodes for which parity with other nodes has not yet been determined is less than or equal to $M$. That is, the number of candidate solutions is less than or equal to $2^{M}$. If $M$ is sufficiently small, e.g., $M\leq 20$, a brute-force search among the candidates is not time-consuming, and hence the best solution ${\bf b}^{*}$ can be quickly obtained. The RQAOA [3] also uses brute-force search.
Example
Figure 4: Schematic illustration of the RQRAO procedure. (a) Initial graph whose edge weights are $+1$ (red) or $-1$ (blue). (b) Three samples of edge energy, $\mathcal{E}^{(1)}_{jk}$, $\mathcal{E}^{(2)}_{jk}$, and $\mathcal{E}^{(3)}_{jk}$, are evaluated using optimized states whose corresponding Hamiltonians are $H^{(1)}_{3}$, $H^{(2)}_{3}$, and $H^{(3)}_{3}$, respectively. Node labels are assigned 1-local Paulis. (c) Ensemble edge energy $\mathfrak{E}_{jk}$ using the three samples illustrated in (b). Closed loop $0-4-2-8-0$ is inconsistent because $(b_{0}=b_{4})\land(b_{4}\neq b_{2})\land(b_{2}=b_{8})\land(b_{8}=b_{0})$ cannot be satisfied. (d) Maximum spanning tree with $|\mathfrak{E}_{jk}|$ edge weights. Red and blue indicate positive and negative edges, respectively. Edge $(0,4)$ is ignored at this step. (e) Modified graph after all leaf nodes of (d) have been removed and the edges in the tree have been reconnected to the remaining nodes. RQRAO repeats steps (a) to (e) until the number of nodes becomes smaller than $M$.
The whole RQRAO procedure is schematically shown in Fig. 4. The problem graph $G$ is shown in Fig. 4, where the red and blue edges have weights $+1$ and $-1$, respectively. The random Pauli assignments and optimization are carried out to obtain the edge energy for each trial $t$, $\mathcal{E}^{(t)}_{jk}$ (Fig. 4. The ensemble edge energy $\mathfrak{E}_{jk}$ is then calculated using $\{\mathcal{E}^{(t)}_{jk}\}^{N}_{t=1}$ (Fig. 4. The inconsistent edges are removed by solving the maximum spanning tree problem, where $|\mathfrak{E}_{jk}|$ are regarded as the edge weights. The rooted tree is constructed by giving the direction from root to leaf, where the root node has been randomly selected (Fig. 4. The edge parities in the rooted tree are determined from leaf to root according to the sign of corresponding $\mathfrak{E}_{jk}$. At the same time, the graph is modified to satisfy the determined parity. Figure 4 represents the modified graph after all edge parities in the rooted tree have been determined. The protocol from Fig. 4 to 4 is repeated until the number of nodes of the modified graph becomes smaller than $M$.
3.2 Theoretical Analysis
The proposed method, RQRAO, combines the RQAOA [3] and QRAO [11]. These existing methods work well experimentally, but there are some theoretical ambiguities: (i) Why does the recursive approach improve the cut weight performance in the RQAOA? (ii) What is the meaning of maximizing the expectation value of the QRAC Hamiltonian, and what is the theoretical background of Pauli rounding in QRAO? To answer these questions, we prove the following theorem:
Theorem 1
The objective of the RQAOA, QRAO, and RQRAO can be unifiedly expressed as the problem of solving
$\rho^{*}=\mathop{\rm argmax}_{\rho}\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\Big{[}\texttt{CW}({\bf b})\Big{]},$ (18)
${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}\in\{0,1\}^{|V|}}\mathcal{P}_{m}({\bf b};\rho^{*}),$ (19)
where $\mathcal{P}_{m}({\bf b};\rho)$ is the probability of obtaining ${\bf b}$ by $(m,1)$-random magic measurements and is proportional to the fidelity between $\rho$ and the $(m,1)$-magic state embedding $\mu({\bf b})$.
From Theorem 1, it is clarified that both the RQAOA and QRAO, let alone RQRAO, are models searching for the highest probability sample of the probability distribution that maximizes the expectation value of the cut weight through the quantum state as the hidden variable. It will be proved later that maximizing the expectation value of the QRAC Hamiltonian, let alone that of the Ising-type Hamiltonian, can be seen as carrying out Eq. (18). In addition, it will also be proved later that both the recursive approach in the RQAOA and Pauli rounding in the QRAO are just methods that approximate methods of Eq. (19).
Of course, Theorem 1 does not say the MAX-CUT problem is efficiently solvable because both maximizing the expected cut weight and obtaining the highest probability sample is, in general, intractable when ${\bf b}$ lives in an extremely high dimensional space, $\{0,1\}^{|V|}$ for $|V|\gg 1$. However, we can say that it is natural to take the ${\bf b}$ with the highest probability because, after solving Eq. (18), which maximizes the expected cut weight, the probability for larger $\texttt{CW}({\bf b})$ is expected to be large. It is therefore reasonable calculate Eqs. (18) and (19) to obtain a better cut weight.
In Sec. 3.2.1, Eq. (18) is proved. Then, Eq. (19) is proved in Sec. 3.2.2. All the proofs of the following theorems and corollaries are deferred to Supplementary Information C.
3.2.1 Proof of Eq. (18)
The objectives of the RQAOA, QRAO, and RQRAO are to maximize ${\rm tr}(H_{m}\rho)$, where $m=1$ for the RQAOA whereas $m\in\{1,2,3\}$ for QRAO and RQRAO. Hence, to prove Eq. (18), we need to show that maximizing ${\rm tr}(H_{m}\rho)$ is equivalent to maximizing $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}$. To show this, let us start with the following theorem and corollary, which are proved for the first time in this paper.
Theorem 2
Let $P^{\otimes k}:=\prod^{k}_{i=1}P_{q^{(i)}}\in{\mathbb{C}}^{2^{n}\times 2^{n}}$ be the $k$-local Pauli observable, where $P_{q}$ represents $1$-local Pauli $P\in\mathfrak{P}\subset\{X,Y,Z\}$ acting on the $q$th qubit, and $q^{(i)}$ is the $i$th qubit index, where $q^{(i)}\neq q^{(j)}$ for $i\neq j$. Suppose the case $|\mathfrak{P}|=m$. Let $\rho\in{\mathbb{C}}^{2^{n}\times 2^{n}}$ be the arbitrary $n$ qubit state, $\mu_{m}({\bf b})\in{\mathbb{C}}^{2^{n}\times 2^{n}}$ be the $n$ qubit $(m,1)$-magic state defined in Eq. (12), in which ${\bf b}\in\{0,1\}^{mn}$ is embedded, and $f_{m}({\bf b},\rho):={\rm tr}(\mu_{m}({\bf b})\rho)$ be the fidelity between $\mu_{m}({\bf b})$ and $\rho$. The following relationship holds:
${\rm tr}\big{(}P^{\otimes k}\rho\big{)}=m^{k/2}\,\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\Big{[}(-1)^{\sum^{k}_{i=1}b^{[q^{(i)}]}_{P}}\Big{]},$ (20)
where $b^{[q]}_{P}\in\{0,1\}$ is a binary mapped to the Pauli $P$ of the $q$th qubit and
$\mathcal{P}_{m}({\bf b};\rho):=\frac{f_{m}({\bf b},\rho)}{\sum_{{\bf b}^{\prime}\in\{0,1\}^{mn}}f_{m}({\bf b}^{\prime},\rho)}.$ (21)
Note that ${\bf b}$ following the probability $\mathcal{P}_{m}$ is the outcome of the $(m,1)$-random magic measurements, whose details are described in Supplementary Information A.
Corollary 1
Let $\rho\in{\mathbb{C}}^{2^{n}\times 2^{n}}$ be the arbitrary $n$ qubit state and $H_{m}$ be the $(m,1)$-QRAC Hamiltonian defined as Eq. (10). The following relationship holds:
${{\rm tr}(H_{m}\,\rho)}$
${=\sum_{(j,k)\in E}\frac{1-m^{2}\,\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}(-1)^{b_{j}+b_{k}}\big{]}}{2}w_{jk}}$
$=m^{2}\,\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}-\frac{m^{2}-1}{2}\sum_{(j,k)\in E}w_{jk},$ (22)
where $\mathcal{P}_{m}$ is defined as Eq. (21).
According to Eq. (22), the expectation value of the $(m,1)$-QRAC Hamiltonian is proportional to the expectation value of the cut weight, where the expectation is taken by the bit strings obtained by $(m,1)$-random magic measurements. Hence, maximizing the expectation value of the $(m,1)$-QRAC Hamiltonian can be regarded as maximizing the expected cut weight. Note that the lower bounds of $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}$ for $w_{jk}=1$ when ${\rm tr}(H_{m}\rho)\geq\texttt{CW}({\bf b}^{*})$ have been given in [11, Theorem 2 and Supplementary Information VII]. They are $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}\geq\frac{5}{9}\texttt{CW}({\bf b}^{*})$ for $m=3$ and $\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}\geq\frac{5}{8}\texttt{CW}({\bf b}^{*})$ for $m=2$. These bounds are reproduced by substituting ${\rm tr}(H_{m}\rho)\geq\texttt{CW}({\bf b}^{*})$ and $\sum_{(j,k)\in E}w_{jk}\geq\texttt{CW}({\bf b}^{*})$ into Eq. (22).
Note that not only the expectation value but also the variance of the cut weight can be obtained using the $(m,1)$-QRAC Hamiltonian as follows.
Corollary 2
Let $\rho\in{\mathbb{C}}^{2^{n}\times 2^{n}}$ be the arbitrary $n$ qubit state and $H_{m}$ be the $(m,1)$-QRAC Hamiltonian defined as Eq. (10). The variance of the cut weight is
${{\rm Var}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}}$
$=\frac{1}{m^{4}}\Big{(}{\rm tr}(H_{m}^{2}\rho)-{\rm tr}(H_{m}\rho)^{2}\Big{)},$ (23)
where $\mathcal{P}_{m}$ is defined as Eq. (21).
3.2.2 Proof of Eq. (19)
First, Eq. (19) is rewritten in the form of the conditional distribution to make it easier to understand. A joint distribution $\mathcal{P}_{m}({\bf b};\rho^{*})$ can be written as the product of the conditional distribution as
$\mathcal{P}_{m}({\bf b};\rho^{*})=\prod^{|V|}_{a=1}\mathcal{P}_{m}(b_{j_{a}}|{\bf b}_{<j_{a-1}};\rho^{*}),$ (24)
where ${\bf b}_{<j_{a-1}}:=\{b_{j_{1}},\ldots,b_{j_{a-1}}\}$. Using this relationship, Eq. (19) can be written as
${b^{\rm R}_{j_{a}}=\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\mathcal{P}_{m}(b_{j_{a}}|{\bf b}^{\rm R}_{<j_{a}};\rho^{*})}$
${\rm for}\ \ \ a=1,2,\ldots,|V|.$ (25)
Because the cut weight is invariant to the bit flip, we can fix one of the $b^{\rm R}_{j}$ to 00 or $1$. In this case, there exists a one-to-one relationship between $\mathcal{P}_{m}({\bf b};\rho^{*})$ and
${\mathcal{P}}_{m}(\mathfrak{p};\rho^{*},T):=\prod^{|V|-1}_{b=1}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}_{<e_{b}};\rho^{*},T),$ (26)
where
$\mathfrak{p}_{e=(c,p)}:=\left\{\begin{array}[]{ll}+&(b_{c}=b_{p})\\[5.69054pt] -&(b_{c}\neq b_{p})\end{array}\right.$ (27)
is a random variable of an edge parity, where “$+$” and “--” indicate positive and negative parities, respectively, $T=\{e_{j_{b}}=(c_{j_{b}},p_{j_{b}})\}^{|V|-1}_{b=1}$ is an arbitrary spanning tree of given graph $G$, and $\mathfrak{p}_{<e_{b}}:=\{\mathfrak{p}_{e_{1}},\ldots,\mathfrak{p}_{e_{b-1}}\}$. Here, the reason for considering the probability of the edge parity on the tree is to exclude inconsistent parities such as a triangle with negative parities. Excluding such an inconsistency does not lose the generality because all possible configurations of solutions can be treated by considering only the probability on the tree. From the view of edge parity, Eq. (19) can be rewritten as
${\mathfrak{p}^{\rm R}_{e_{b}}=\mathop{\rm argmax}_{\mathfrak{p}_{e_{b}}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}}|\mathfrak{p}^{\rm R}_{<e_{b}};\rho^{*},T^{*})}$
${\rm for}\ \ \ b=1,2,\ldots,|V|-1,$ (28)
where
$T^{*}:=\mathop{\rm argmax}_{T}\max_{\mathfrak{p}\in\{+,-\}^{|V|-1}}{\mathcal{P}}_{m}(\mathfrak{p};\rho^{*},T).$ (29)
In the following, we show that QRAO uses Eq. (25), whereas both the RQAOA and RQRAO use Eq. (28) to obtain the solution. All of them assume a moderate approximation on a probability distribution. The following corollary is frequently used for the proof.
Corollary 3
Define $\mathcal{E}_{j}:={\rm tr}(P_{\langle j\rangle}\,\rho)$ and $\mathcal{E}_{jk}:={\rm tr}(P_{\langle j\rangle}P_{\langle k\rangle}\,\rho)$. When the bit string ${\bf b}=b_{|V|-1}\ldots b_{1}b_{0}$ is obtained by the $(m,1)$-random magic measurements, ${\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)$, the probabilities of $b_{j}=0$ and $b_{j}=1$ are
$\begin{array}[]{lll}\mathcal{P}(b_{j}=0)&=&\displaystyle{\frac{1}{2}+\frac{1}{2\sqrt{m}}\mathcal{E}_{j}}\\[14.22636pt] \mathcal{P}(b_{j}=1)&=&\displaystyle{\frac{1}{2}-\frac{1}{2\sqrt{m}}\mathcal{E}_{j}}\end{array},$ (30)
respectively, whereas the probability of $b_{j}=b_{k}$ and $b_{j}\neq b_{k}$ are
$\begin{array}[]{lll}\mathcal{P}(b_{j}=b_{k})&=&\left\{\begin{array}[]{ll}\displaystyle{\frac{1}{2}+\frac{1}{2m}\mathcal{E}_{j}\mathcal{E}_{k}}&(q^{(j)}=q^{(k)})\\[14.22636pt] \displaystyle{\frac{1}{2}+\frac{1}{2m}\mathcal{E}_{jk}}&(q^{(j)}\neq q^{(k)})\end{array}\right.\\[28.45274pt] \mathcal{P}(b_{j}\neq b_{k})&=&\left\{\begin{array}[]{ll}\displaystyle{\frac{1}{2}-\frac{1}{2m}\mathcal{E}_{j}\mathcal{E}_{k}}&(q^{(j)}=q^{(k)})\\[14.22636pt] \displaystyle{\frac{1}{2}-\frac{1}{2m}\mathcal{E}_{jk}}&(q^{(j)}\neq q^{(k)})\end{array}\right.\end{array},$ (31)
respectively, where $q^{(j)}$ is a qubit with $b_{j}$ embedded.
Equation (30) is consistent with the upper bound of the decoding probability, $\frac{1}{2}+\frac{1}{2\sqrt{m}}$ [1], which is recovered when $|\mathcal{E}_{j}|=1$.
Pauli rounding
Assume all $b_{j_{a}}$ for $a=1,2,\ldots,|V|$ are independent. Under this assumption, Eq. (24) is approximated as
${\mathcal{P}}^{\rm Pauli}_{m}({\bf b};\rho^{*}):=\prod^{|V|}_{a=1}\mathcal{P}_{m}(b_{j_{a}};\rho^{*}).$ (32)
Substituting Eq. (32) into Eq. (25) gives
$\begin{array}[]{lll}b^{\rm Pauli}_{j_{a}}&:=&\displaystyle{\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\mathcal{P}_{m}(b_{j_{a}};\rho^{*})}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{b_{j_{a}}\in\{0,1\}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{j_{a}}}}{2\sqrt{m}}\mathcal{E}_{j_{a}}\Bigg{)}}\\[14.22636pt] &=&\left\{\begin{array}[]{ll}0&(\mathcal{E}_{j_{a}}>0)\\ 1&(\mathcal{E}_{j_{a}}<0)\end{array}\right.,\end{array}$ (33)
where the second equality uses Corollary 3. Interestingly, ${\bf b}^{\rm Pauli}$ is equivalent to the result of the Pauli rounding of QRAO [11]. This comes out that the Pauli rounding seeks the variable with the highest probability of ${\mathcal{P}}_{m}({\bf b};\rho)$ under the assumption of Eq. (32).
Tree rounding
Assume all $\mathfrak{p}_{e_{b}}$ for $b=1,2,\ldots,|V|-1$ are independent. Under this assumption, Eq. (26) is approximated as
${\mathcal{P}}^{\rm tree}_{m}(\mathfrak{p};\rho^{*},T):=\displaystyle{\prod_{b=1}^{|V|-1}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}};\rho^{*},T)}.\$ (34)
Substituting Eq. (34) into Eq. (28) gives
$\begin{array}[]{lll}\mathfrak{p}^{\rm tree}_{e_{b}=(c,p)}&:=&\displaystyle{\mathop{\rm argmax}_{\mathfrak{p}_{e_{b}}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}};\rho^{*},T^{*})}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{(b_{c},b_{p})\in\{0,1\}^{2}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{c}+b_{p}}}{2m}\mathcal{E}_{cp}\Bigg{)}}\\[14.22636pt] &=&\left\{\begin{array}[]{ll}+&(\mathcal{E}_{cp}>0)\\ -&(\mathcal{E}_{cp}<0)\end{array}\right.,\end{array}$ (35)
where the second equality uses Corollary 3, and
$\begin{array}[]{lll}T^{*}&=&\displaystyle{\mathop{\rm argmax}_{T}\max_{\mathfrak{p}}{\mathcal{P}}^{\rm tree}_{m}(\mathfrak{p};\rho^{*},T)}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{T\ni\{(c,p)\}}\max_{(b_{c},b_{p})\in\{0,1\}^{2}}\Bigg{(}\frac{1}{2}+\frac{(-1)^{b_{c}+b_{p}}}{2m}\mathcal{E}_{cp}\Bigg{)}}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{T\ni\{(c,p)\}}\big{|}\mathcal{E}_{cp}\big{|}}\end{array}$ (36)
is a maximum spanning tree. Computing Eq. (35) can be carried out by a two-step calculation. In the first step, the maximum spanning tree $T^{*}$ is computed with the edge weights $|\mathcal{E}_{cp}|$ for edge $(c,p)$. In the second step, the edge parity is determined by the sign of corresponding $\mathcal{E}_{cp}$. We refer to this rounding method as tree rounding.
RQRAO with $N=1$, which has no recursive steps, is equivalent to using tree rounding. Because both Pauli and tree rounding assume that all random variables are independent, it is expected that QRAO and RQRAO ($N=1$) will yield a similar cut weight.
Recursive rounding
Both Pauli and tree rounding impose the strong approximation on $\mathcal{P}_{m}$ that all random variables are independent. This is because conditional sampling is in general intractable to carry out. For this reason, let us impose the condition $\mathfrak{p}_{<e_{b}}=\mathfrak{p}^{*}_{<e_{b}}$ on the optimization problem as the constraint instead of on $\mathcal{P}_{m}$. Of course, these two approaches are not exactly equivalent, but imposing the condition as the constraint of the optimization should be a good approximation to the conditional sampling. The specific steps are given below.
First, one parity of $e_{1}$ is determined using the optimization result, $\rho^{*}_{1}$ as
$\mathfrak{p}^{*}_{e_{1}}=\mathop{\rm argmax}_{\mathfrak{p}_{e_{1}}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e_{1}};\rho^{*}_{1}),$ (37)
where
$e_{1}=\mathop{\rm argmax}_{e}\max_{\mathfrak{p}_{e}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e};\rho^{*}_{1}).$ (38)
According to Corollary 3, this can be rewritten as
$\mathfrak{p}^{\rm rec}_{e_{1}=(c_{1},p_{1})}=\left\{\begin{array}[]{ll}+&(\mathcal{E}_{c_{1}p_{1}}>0)\\ -&(\mathcal{E}_{c_{1}p_{1}}<0)\end{array}\right.,$ (39)
where
$\begin{array}[]{lll}(c_{1},p_{1})&=&\displaystyle{\mathop{\rm argmax}_{(c,p)\in E}\Bigg{(}\frac{1}{2}+\frac{1}{2m}\big{|}\mathcal{E}_{cp}\big{|}\Bigg{)}}\\[14.22636pt] &=&\displaystyle{\mathop{\rm argmax}_{(c,p)\in E}\big{|}\mathcal{E}_{cp}\big{|}}.\end{array}$ (40)
After that, the given graph $G$ is modified to satisfy $\mathfrak{p}_{e_{1}}=\mathfrak{p}^{*}_{e_{1}}$, which refers to $G_{2}$. Next, the second optimization is carried out with the $(m,1)$-QRAC Hamiltonian based on $G_{2}$, and the optimized state $\rho^{*}_{2}$ is obtained. Because the modified graph is forced to be $\mathfrak{p}_{e_{1}}=\mathfrak{p}^{*}_{e_{1}}$, it can be seen that $\rho^{*}_{2}$ is conditioned on $\mathfrak{p}_{e_{1}}=\mathfrak{p}^{*}_{e_{1}}$. Hence, the determined parity of the second step can be written as
$\mathfrak{p}^{*}_{e_{2}}=\mathop{\rm argmax}_{\mathfrak{p}_{e_{2}}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e_{2}};\rho^{*}_{2}(\mathfrak{p}_{e_{1}}=\mathfrak{p}^{*}_{e_{1}}))$ (41)
where
$e_{2}=\mathop{\rm argmax}_{e}\max_{\mathfrak{p}_{e}\in\{+,-\}}\mathcal{P}_{m}(\mathfrak{p}_{e};\rho^{*}_{2}(\mathfrak{p}_{e_{1}}=\mathfrak{p}^{*}_{e_{1}})).$ (42)
Of course, this could also be rewritten using the edge energy $\mathcal{E}_{cp}$, but we will leave that out as it is self-explanatory. Repeating this procedure gives
${\mathfrak{p}^{*}_{e_{b}}=\mathop{\rm argmax}_{\mathfrak{p}_{e_{b}}}\mathcal{P}_{m}(\mathfrak{p}_{e_{b}};\rho^{*}_{b}(\mathfrak{p}_{<e_{b}}=\mathfrak{p}^{*}_{<e_{b}}))}$
$\ \ \ {\rm for}\ \ \ b=1,\ldots,|V|-1.$ (43)
Finally, $\mathfrak{p}^{*}=\{\mathfrak{p}^{*}_{1},\ldots,\mathfrak{p}^{*}_{|V|-1}\}$ is obtained. The corresponding bit string ${\bf b}^{\rm rec}$ can be easily recovered from $\mathfrak{p}^{*}$ by fixing one of the $b^{\rm rec}_{j}$ to 00 or $1$. The condition $\mathfrak{p}_{<e_{b}}=\mathfrak{p}^{*}_{<e_{b}}$ is indirectly imposed through $\rho^{*}_{b}$ to $\mathcal{P}_{m}$, and hence this rounding method imposes an approximation that is weaker than that of Pauli and tree rounding. We refer to this rounding method as recursive rounding.
In recursive rounding, the edge that determines the parity is selected using the absolute value of its edge energy (Eq. (42)), and its parity is determined by the sign of the corresponding edge energy (Eq. (41)). This process is the same as explained in Sec. 2.2 and hence, it can be said that the RQAOA adopts recursive rounding. The main idea behind RQAOA is to approximate $\mathop{\rm max}_{{\bf b}\in\{0,1\}^{|V|}}{\rm tr}(H_{1}|{\bf b}\rangle\langle{\bf b}|)$ [3], but in actuality, it approximates $\mathop{\rm max}_{{\bf b}\in\{0,1\}^{|V|}}\mathcal{P}_{1}({\bf b};\rho^{*})=\mathop{\rm max}_{{\bf b}\in\{0,1\}^{|V|}}{\rm tr}(\rho^{*}|{\bf b}\rangle\langle{\bf b}|)$.111 Since $\mu_{1}({\bf b})=|{\bf b}\rangle\langle{\bf b}|$, $f_{1}({\bf b},\rho)={\rm tr}(\mu_{1}({\bf b})\rho)$, and $\sum_{{\bf b}\in\{0,1\}^{|V|}}f_{1}({\bf b},\rho)=1$, $\mathcal{P}_{1}({\bf b};\rho)=f_{1}({\bf b},\rho)/\sum_{{\bf b}^{\prime}}f_{1}({\bf b}^{\prime},\rho)={\rm tr}(|{\bf b}\rangle\langle{\bf b}|\rho)$. By contrast, the rounding for RQRAO with $N>1$ can be considered to be an intermediate approach between tree rounding and recursive rounding because the edge parities that are being determined simultaneously in one step are assumed to be independent.
4 Numerical Experiments
In this section, we show the results of the numerical experiments of the MAX-CUT problem using RQRAO and other existing algorithms. The main claim is that RQRAO achieves a better cut weight than existing algorithms and is comparable to the state-of-the-art classical heuristic.
Datasets
We used three graph datasets.
The first dataset is used to show the validity of Theorem 1. That is, it shows that the bit string with the highest probability yields a good cut weight after maximizing the expectation of the $(m,1)$-QRAC Hamiltonian. Only one graph instance is generated by a machine-independent graph generator rudy [19] using -rnd_graph 14 50 0 -random 0 1 0 -times 2 -plus -1, where the number of nodes is $14$, the graph density is $0.5$, and the edge weights are set to ±$\pm 1$ uniformly at random. This instance has only one MAX-CUT solution, ${\bf b}=00001101101010$, except for its bit-flip solution, where the maximum cut weight is 12. We call this graph instance Rnd14.
The second dataset is the benchmark graph dataset Gset [15] (http://www.stanford.edu/~yyye/yyye/Gset/). Gset includes random, toroidal, and almost planar graphs with $|V|=800$ to $20000$ indexed by G1 to G81, which are generated by rudy. Only G1, G6, G14, and G18 are used for evaluation, all of which have 800 nodes. G1 and G6 are random graphs, G11 is a toroidal graph, and G14 and G18 are the combined graphs of two planar graphs. Graphs G1 and G6, and graphs G14 and G18 have the same edge structure except for the weights, where G1 and G14 have $+1$ edge weights whereas G6 and G18 have edge weights of ±$\pm 1$ assigned at random. The edge weights of G11 are ±$\pm 1$ assigned at random. We call this graph dataset Gset800.
The third dataset consists of 3-regular graphs with ±$\pm 1$ edge weights generated at random. The number of nodes varies from $10$ to $10^{4}$. Ten graphs are generated for each number of nodes. These graphs are used to evaluate the scalability with respect to the cut weight and the runtime of the algorithms. We call this graph dataset Rnd3R.
Baselines
We compared our model with GW [7], CirCut [16], QRAO [11], and RQAOA [3]. GW is a well-known classical approximation algorithm based on semidefinite programming. CirCut is the Fortran program name of the rank-two relaxation algorithm that is a classical heuristic algorithm and empirically outperforms other classical heuristics on various graph instances [20]. The details of GW and CirCut are described in Sec. 5.1. The RQAOA is a QAOA-based algorithm whose details are explained in Sec. 2.2. QRAO is a quantum relaxation-based optimization method whose details are explained in Sec. 2.3.
Implementation Details
The matrix product state (MPS) ansatz is used for $|\psi({\boldsymbol{\theta}})\rangle$, where all the matrices in the MPS are treated as trainable parameters (see Supplementary Information D for details). The default hyperparameters for RQRAO used in the numerical experiments are listed in Table 1.
We implemented our model in Python, and the MPS optimization is implemented using PyTorch version 1.10.0 [21]. For the optimizer, we used L-BFGS [22] with line search using strong Wolfe conditions. The termination tolerance for changing the function value and parameters is set to $10^{-2}$. Implementation details for GW, CirCut, the RQAOA, and QRAO are described in Supplementary Information E.
All experiments were carried out on a single Intel(R) Core(TM) i9-9900X CPU @ 3.50GHz with 32GB (8GBx4) DDR4-2666 Quad-Channel.
Table 1: Default hyperparameters for RQRAO.
Hyperparameter Symbol Value
# embeddings / qubits $m$ $3$
# ensemble $N$ $20$
Scale factor $S$ $2$
Bond dimension $\chi$ $2$
Brute-force search threshold $M$ $10$
Details of the Experiments
Using the Rnd14 dataset, Theorem 1 was verified. First, $(m,1)$-QRAC Hamiltonians were randomly generated 100 times each for $m=1,2,$ and $3$. For each Hamiltonian, the eigenstate with the maximum eigenvalue and optimized state with bond dimension $\chi=2$ were then calculated. Using each state, the cut weight with the highest probability was calculated. We call this experiment ExpVer.
Using the Gset800 dataset, the procedure was run 10 times on the same graph instance and the best cut of the trials was recorded. RQRAO and all baseline methods were used. We call this experiment ExpGset.
Using the Rnd3R dataset, the procedure was run once for each graph instance and the cut weight and elapsed time were recorded. The cut weight is normalized with respect to that of the GW algorithm and is referred to as the relative cut weight. Using Rnd3R, two experiments were carried out. The first one evaluated the scalability. RQRAO and all baseline methods were used. We call this experiment ExpScale. The second one analyzed the sensitivity of the RQRAO hyperparameters. Only one hyperparameter of RQRAO was varied from its default value (Table 1). We call this experiment ExpHyp.
4.1 Results and Discussion
Verification of Theorem 1 (ExpVer)
(a) MaxEig
(b) $\chi=2$
Figure 5: Solutions with the highest probability of 100 trials for the same graph instance, Rnd14, where the state used was (a) the eigenstate corresponding to the maximum eigenvalue and (b) the state obtained through optimization with $\chi=2$.
Figure 5 summarizes the results of ExpVer. If one can obtain the eigenstate with the maximum eigenvalue, $(1,1)$-QRAC is the best choice because the cut weight with the highest probability for $(1,1)$-QRAC is always the MAX-CUT solution (Fig. 5(a)). By contrast, in the case of $(2,1)$ and $(3,1)$-QRACs, the solution with the highest probability is sometimes not the MAX-CUT solution. This occurs when low cut weight solutions are embedded near the MAX-CUT solution, and higher but not the maximum cut weight solutions are aggregated somewhere else in Hilbert space. In such a case, the state with the maximum eigenvalue is placed in its aggregation, which could be far from the MAX-CUT solution. This never occurs in the case of $(1,1)$-QRAC, because all embeddings are orthogonal to each other. There are drawbacks due to these accidental cases. However, $(2,1)$ and $(3,1)$-QRACs still have relatively higher cut weights. Therefore, it makes sense to choose a sample with the highest probability as a solution to obtain a higher cut weight solution, which is what the RQAOA, QRAO, and RQRAO attempt to implement according to Theorem 1, if the state corresponding to the maximum eigenvalue can be obtained. However, in practice, the state obtained through optimization is not the eigenstate with the maximum eigenvalue. In this case, the cut weight with the highest probability for both $(2,1)$- and $(3,1)$-QRACs is much better than that for $(1,1)$-QRAC (Fig. 5(b)). This suggests that $(2,1)$- or $(3,1)$-QRAC is a better choice for obtaining a higher cut weight solution when the state is obtained through optimization. One reason for this difference comes from the difficulty of optimization. Optimization using the $(2,1)$- and $(3,1)$-QRAC Hamiltonians would be easier than using the $(1,1)$-QRAC because the numbers of qubits of $(2,1)$- and $(3,1)$-QRAC Hamiltonians are less than that of $(1,1)$-QRAC. The other reason is the difference in the landscape of the objective functions. However, the details have not yet been clarified and are left for future work.
Originally, the MAX-CUT problem in an Ising-type formulation was to find the eigenstate with the maximum eigenvalue, which is known to be QMA-hard [12]. Remember that the variable with the highest probability is embedded in the magic state with the highest fidelity to the current quantum state. This suggests that the MAX-CUT problem has now been changed into a problem of finding the magic state with the highest fidelity to the optimized quantum state. Of course, finding the magic state with the largest fidelity is also intractable, and the largest fidelity sample is not always the MAX-CUT solution, as seen in Fig. 5(a). However, in the practical case, we can expect that a better solution can be obtained using $(2,1)$- and $(3,1)$-QRACs rather than using $(1,1)$-QRAC, as shown in Fig. 5(b). In other words, the MAX-CUT problem could be successfully relaxed by the QRAC Hamiltonian.
Gset (ExpGset)
Table 2: Comparison with existing methods on MAX-CUT problems from the Gset. All instances have $|V|=800$. All best-known values were reported in [23] with the breakout local search algorithm. GW: Goemans–Williamson algorithm [7], RQAOA: recursive quantum approximate optimization algorithm [3], QRAO: quantum random access optimization [11], RQRAO: recursive quantum random access optimization (proposed).
Instance Best-known GW CirCut QRAO RQAOA ${\rm RQRAO}_{N=1}$ ${\rm RQRAO}_{N=20}$
G1 11624 11467 11622 11453 11441 11466 11562
G6 2178 2013 2178 1980 1862 1980 2148
G11 564 536 556 540 562 544 564
G14 3064 2999 3052 2971 2965 2965 3043
G18 992 924 987 929 882 920 980
The cut weight on the Gset dataset for each algorithm is summarized in Table 2. RQRAO with $N=20$ exceeds both the classical approximation algorithm (GW) and quantum algorithms (QRAO and the RQAOA), but is slightly worse than the classical heuristic algorithm (CirCut) on all graphs except for G11.
As expected in Sec. 3.2.2, the cut weight of RQRAO with $N=1$ is very similar to that of QRAO because both assume all random variables are independent. The cut weight performance of RQRAO increases when the number of ensembles $N$ is increased, which confirms the effectiveness of using an ensemble, which is heuristically introduced in our method.
Scalability (ExpScale)
Figure 6: Comparison of scalability with existing methods. Ten 3-regular graphs with the same number of nodes were randomly generated whose edge weights were ±$\pm 1$ at random. The markers indicate the means. (a) Relative cut weight with respect to the cut weight obtained by the GW algorithm; (b) Algorithm run time. The gray dashed line is a fitted curve using $({\rm Elapsed\ time\ [sec]})=a({\rm Number\ of\ nodes})^{b}$
Figure 6: Comparison of scalability with existing methods. Ten 3-regular graphs with the same number of nodes were randomly generated whose edge weights were ±$\pm 1$ at random. The markers indicate the means. (a) Relative cut weight with respect to the cut weight obtained by the GW algorithm; (b) Algorithm run time. The gray dashed line is a fitted curve using $({\rm Elapsed\ time\ [sec]})=a({\rm Number\ of\ nodes})^{b}$
Figure 6 summarizes the results on $3$-regular graphs. Although the relative cuts of QRAO and the RQAOA are lower than 1, that of RQRAO exceeds 1 when the number of nodes is larger than 200 and is larger than that of CirCut (Fig. 6). In addition, although the absolute runtime of RQRAO is longer than that of CirCut, the order of empirical time complexity of these methods is close (Fig. 6). To quantitatively compare the empirical time complexity of the methods, the runtime curves were fitted using
${({\rm Wall\ time\ [sec]})=a({\rm Number\ of\ nodes})^{b}}$
using only data with a large number of nodes. The fitted values of $b$ are $1.34$ for CirCut, $2.78$ for the RQAOA, $1.73$ for QRAO, $1.44$ for RQRAO, and $2.44$ for GW. That is, the empirical time complexity of the proposed method is much lower than that of GW and the RQAOA. Moreover, it is comparable to that of CirCut. Because most of the computational time is spent on optimizing the variational states, the computational time of the proposed method can be reduced by improving the optimization algorithm. For example, the runtime cost of RQRAO can be reduced by parallel computation in the ensemble part. Further reductions in the computational costs of RQRAO are future work.
Sensitivity Analysis (ExpHyp)
Figure 7: Results of the ablation study for the number of embeddings per qubit $m$. The problem graph instances are $3$-regular graphs with ±$\pm 1$ edge weights. For each number of nodes, $10$ graph instances were randomly generated.
Only the results for varying the number of embeddings per qubit $m$ are described here. Other hyperparameter sensitivities are summarized in Supplemental Information F. In Fig. 7, we can see the clear advantage of using $m>1$ over $m=1$, which is that only the cut weight in the cases of $m>1$ overcomes that of GW. When $m=1$, the optimized state may be stacked into a bad local minimum, as shown in Fig. 5(b), and this is the reason for the lower cut weight. Almost no runtime dependency on $m$ is seen in Fig. 7, which means there is no disadvantage to using $m>1$ as long as the given graph is sufficiently sparse and the MPS ansatz is adopted.
5 Related Work
5.1 Classical Algorithm for the MAX-CUT problem
Before reviewing the classical algorithms for the MAX-CUT problem, we reformulate Definition 1 to make the existing classical algorithms easier to understand. Because $(-1)^{b_{j}+b_{k}}=v_{j}v_{k}$ where $v_{j}:=(-1)^{b_{j}}$, the MAX-CUT problem can be reformulated as
$\max_{{\bf X}}\texttt{CW}_{\rm X}({\bf X}),$ (44)
where
$\texttt{CW}_{\rm X}({\bf X}):={\rm tr}({\bf w}{\bf F}({\bf X})),$ (45)
${\bf F}({\bf X}):=({\bf 1}-{\bf X})/2$, and ${\bf X}:={\bf v}{\bf v}^{\mathsf{T}}$. By definition, ${\bf X}$ is restricted as ${\rm diag}({\bf X})={\bf 1}$, ${\rm rank}({\bf X})=1$, and ${\bf X}\succeq 0$.
The best known theoretically guaranteed generic algorithm for the MAX-CUT problem is the GW [7] as stated in Sec. 1. In the GW algorithm, the constraint that ${\rm rank}{(\bf X})=1$ is removed. This transforms the MAX-CUT problem to a semidefinite program (SDP), which can be solved in polynomial time for the number of nodes. After solving the SDP, the solution ${\bf v}\in\{\pm 1\}^{|V|}$ needs to be extracted from the obtained ${\bf X}$. When ${\bf X}$ is written in the form ${\bf X}={\bf vv}^{\mathsf{T}}$, the solution is ${\bf v}$. If not, as is generally the case here, the obtained ${\bf X}$ is decomposed as ${\bf X}={\bf L}{\bf L}^{\mathsf{T}}$ using, for example, Cholesky decomposition, where ${\bf L}=\begin{bmatrix}{\bf l}_{1}&\dots&{\bf l}_{|V|}\end{bmatrix}^{\mathsf{T}}\in\mathbb{R}^{|V|\times|V|}$ is a lower triangular matrix and ${\bf l}_{i}\in{\mathbb{R}}^{|V|}$ is a unit vector. Then, the solution $\{v_{i}\}$ is determined by the random projection of ${\bf l}_{i}$ as $v_{i}=1$ if ${\bf l}_{i}^{\mathsf{T}}{\bf r}\geq 0$ and $v_{i}=-1$ otherwise. Here, ${\bf r}\in{\mathbb{R}}^{|V|}$ is a random vector uniformly distributed on the $|V|$-dimensional unit sphere. This can be viewed as dividing the set $\{{\bf l}_{i}\}^{|V|}_{i=1}$ into two sets by a hyperplane with ${\bf r}$ as the normal vector. When the hyperplane cutting is performed uniformly at random, the lower bound of the expected cut weight is $0.878\,\texttt{CW}({\bf b}^{*})$ [7]. This bound is explained as follows. According to [7, Theorem 3.1], the MAX-CUT problem can be further reformulated as
$\{{\bf l}^{\rm opt}_{i}\}:=\mathop{\rm argmax}_{\{{\bf l}_{i}\}}{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}_{i}\})}[\texttt{CW}({\bf b})]$ (46)
where
$\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}_{i}\}):=\frac{{\rm arccos}({\bf l}_{j}^{\mathsf{T}}{\bf l}_{k})}{\pi}.$ (47)
In the GW algorithm, the objective is rewritten as
$\{{\bf l}^{*}_{i}\}:=\mathop{\rm argmax}_{\{{\bf l}_{i}\}}{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm GW}({\bf b};\{{\bf l}_{i}\})}[\texttt{CW}({\bf b})]$ (48)
where
${\mathcal{P}^{\rm GW}(b_{j}\neq b_{k};\{{\bf l}_{i}\}):=\frac{1-{\bf l}_{j}^{\mathsf{T}}{\bf l}_{k}}{2}}$
$\leq\frac{1}{0.878}\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}_{i}\}).$ (49)
The last inequality of Eq. (49) comes from [7, Lemma 3.4, Lemma 3.5]. 222 $\min_{x\in[-1,1]}\frac{{\rm arccos}(x)}{\pi}\big{(}\frac{1-x}{2}\big{)}^{-1}\approx 0.878$. The right-hand side of Eq. (48) is an SDP, and hence it can be solved in polynomial time. The resulting ${\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})}[\texttt{CW}({\bf b})]$, which is the expected cut weight of random hyperplane cutting, is guaranteed to be greater than $0.878\,{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{\rm opt}_{i}\})}[\texttt{CW}({\bf b})]=0.878\max_{{\bf b}}\texttt{CW}({\bf b})$ because
${\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}^{*}_{i}\})\geq 0.878\,\mathcal{P}^{\rm GW}(b_{j}\neq b_{k};\{{\bf l}^{*}_{i}\})}$
${\geq 0.878\,\mathcal{P}^{\rm GW}(b_{j}\neq b_{k};\{{\bf l}^{\rm opt}_{i}\})}$
$=0.878\,\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}^{\rm opt}_{i}\}).$ (50)
Equation (48) is very similar to Eq. (18), but the rounding process of the GW is not the same as that of Eq. (19). Instead of carrying out ${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}}\mathcal{P}^{\rm GW}({\bf b};\{{\bf l}^{*}_{i}\})$ or ${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}}\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})$, a number of ${\bf b}$ obeying $\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})$ are sampled by random hyperplane cutting and then, the best $\texttt{CW}({\bf b})$ is adopted as the candidate solution.
To solve the GW algorithm faster with less memory, various SDP solvers have been developed [24]. The runtime of the state-of-the-art SDP solvers are $\widetilde{O}(\sqrt{n}(mn^{2}+m^{\omega}+n^{\omega})\log(1/\epsilon)))$ [25] and $\widetilde{O}((\sqrt{n}(m^{2}+n^{4})+m^{\omega}+n^{2\omega})\log(1/\epsilon)$ [26], where $n$ is the problem size, $m$ the number of constraints, $\omega$ the matrix multiplication constant $\leq 2.372927$, and $\epsilon$ the relative accuracy. In the case of the MAX-CUT problem, the runtime is $\widetilde{O}(|V|^{3.5}\log(1/\epsilon))$ because $n=m=|V|$. The overall time complexity for the GW algorithm depends on the runtime of SDP because the naïve Cholesky decomposition takes $O(n^{3})=O(|V|^{3})$ to run, which is faster than the state-of-the-art SDP solver, as long as the number of trials of random hyperplane cutting is less than the time complexity of the SDP. Note that the computational complexity of SDP with respect to the number of nodes can be further reduced to $\widetilde{O}(|V|/\epsilon^{3.5})$ using a specific algorithm for the MAX-CUT , by taking a disadvantage with respect to the accuracy [27].
Although there is no approximation ratio guarantee nor computational complexity bound, there are several classical heuristics that are empirically faster and have higher approximation ratios than the GW algorithm. The rank-two relaxation algorithm [16] relaxed ${\rm rank}({\bf X})=1$ to ${\bf X}={\bf AA}^{\mathsf{T}}$, where ${\bf A}=\begin{bmatrix}{\bf a}_{1}&\dots&{\bf a}_{|V|}\end{bmatrix}^{\mathsf{T}}$, ${\bf a}_{j}=\begin{bmatrix}\cos\theta_{j}&\sin\theta_{j}\end{bmatrix}^{\mathsf{T}}$, and $\theta_{j}\in[0,2\pi)$. Similar to the GW, the objective of the rank-two relaxation can be reformulated as
$\{\theta^{*}_{j}\}:=\mathop{\rm argmax}_{\{\theta_{j}\}}\mathbb{E}_{\mathcal{P}^{\rm R2}({\bf b};\{\theta_{j}\})}[\texttt{CW}({\bf b})],$ (51)
where
$\mathcal{P}^{\rm R2}(b_{j}\neq b_{k};\{\theta_{j}\})=\sin^{2}\Big{(}\frac{\theta_{j}-\theta_{k}}{2}\Big{)}.$ (52)
As with the GW algorithm, the rank-two relaxation does not take the sample with the highest probability as a candidate. Moreover, because the hyperplane is only two-dimensional, random hyperplane cutting is not used. Instead, exhaustive search is used, expressed as
${\bf b}^{\rm R2}=\texttt{CW}({\bf b}^{*}(\alpha^{*})),$ (53)
where
$b^{*}_{j}(\alpha)=\left\{\begin{array}[]{lll}0&{\rm for}&\cos(\theta^{*}_{j}+\alpha)\geq 0\\[5.69054pt] 1&{\rm for}&\cos(\theta^{*}_{j}+\alpha)<0\\ \end{array}\right.$ (54)
and
$\alpha^{*}=\mathop{{\rm argmax}}_{\alpha}\texttt{CW}({\bf b}^{*}(\alpha)).$ (55)
To optimize ${\theta_{j}}$, a gradient-based optimizer is used. In addition, $\{1,2\}$-local search and restarting optimization are incorporated. Here, $k$-local search is an exhaustive method that searches the entire solution within Hamming distance $k$ of the current solution. Breakout local search [23] repeats the 1-local search and perturbation of the current solution using a tabu list. Once a local search falls into a local optimum, perturbation is added to the current solution, and then the local search is restarted. In the perturbation phase, one of three kinds of perturbation is applied with a specific probability. The tabu list is used to prohibit the perturbation of the listed nodes so that the same local minimum will not be visited again.
5.2 Quantum Algorithms for the MAX-CUT problem
The quantum algorithms for the MAX-CUT problem are categorized into two research streams. One is quantum approximate optimization algorithms (QAOAs) [8, 9] and the other is quantum algorithms for SDP [28, 29, 30, 31, 32, 33, 34, 35, 36]. Here, quantum annealing is excluded.
QAOA variants are well summarized in a recent review [10] and are not re-summarized here.
The quantum SDP algorithms are categorized into two types. One encodes the solution to a density matrix and the other uses QRAM [37] to store the solution.
For the density matrix encoding approach, there are two types of encoding methods. One uses a Gibbs state to represent a solution [28, 29, 30, 31, 33], which can be regarded as a solution that is encoded in the corresponding Hamiltonian. The Hamiltonian used in the Gibbs state is updated by the matrix multiplicative weight update method [38] or the matrix exponentiated gradient update method [39]. The other density matrix encoding approach uses a parametrized quantum circuit to represent a solution [34, 35, 36]. The parameter is optimized by frameworks using the variational quantum algorithm [40, 41].
The state-of-the-art quantum SDP algorithm [42] is based on the primal-dual central path method, which uses quantum linear algebra and stores the solution in QRAM. The runtime of this quantum SDP algorithm is $O((mn^{1.5}+n^{3}){\rm poly}(\kappa,\log(mn/\epsilon))$ where $\kappa$ is the condition number of the matrices appearing in the algorithm, and $\epsilon$ is the accuracy parameter. In the case of the MAX-CUT problem, we have $\widetilde{O}(|V|^{3}{\rm poly}(\kappa,\log(|V|^{2}/\epsilon))$, which is a slight improvement on its classical counterpart, which is $\widetilde{O}(|V|^{3.5}\log(1/\epsilon))$.
Although these algorithms have the potential to solve an SDP faster than classical ones, the approximation ratio is the same as that of the GW algorithm, which is empirically worse than that of classical heuristics.
Very recently, [43] proposed the quantum-inspired algorithm based on the observation that ADAPT-QAOA [44] can be approximately expressed as the Clifford circuit, which can be efficiently simulatable using a classical computer. Their model showed the comparable cut weight performance to the GW with several thousand hyperplane cuttings.
5.3 QRACs
The concept that encodes $m$ bits into $n$ (qu)bits was originally proposed by Wiesner as conjugate coding [45] and later rediscovered by Ambainis et al. [1] as (quantum) random access codes ((Q)RACs). $(m,n,p)$ random access encoding is defined as the function that maps $m$ classical bits into $n$ (qu)bits with a decoding probability at least $p$ [1, Definition 1.1]. The minimum number of $n$ with $p>1/2$ is bounded as $(1-\eta(p))m$ for both classical [1, Theorem 2.1] and quantum [46, Theorem 2.3] settings, where $\eta(p)$ is the binary entropy function. In the case of $(2,1,p)$- and $(3,1,p)$-QRACs, there exists $p=\frac{1}{2}+\frac{1}{2\sqrt{2}}$ [1, Lemma 3.1] and $p=\frac{1}{2}+\frac{1}{2\sqrt{3}}$ [1, attributed to Chuang], respectively, which are consistent with our results (Eq. (30)) in the case of $\mathcal{E}_{j}\in\{\pm 1\}$. It was proved that $(2^{n},n,>1/2)$ classical random access encoding [47, Theorem 5.5] and $(2^{2n},n,>1/2)$ quantum random access encoding [48, Theorem 2] do not exist. From these observations, $(3,1)$-QRAC shows the maximum number of classical bits that can be embedded in one qubit with a decoding probability that is larger than $1/2$. This is why we only consider $m\in\{1,2,3\}$ in this study. In addition to $(m,1,p)$-QRAC, $(m,2,p)$-QRAC is particularly of interest and well studied in terms of the theoretical bound [48, 49, 50, 51] and concrete construction scheme [52]. For arbitrary $(m,n,p)$-QRAC, the upper bound of decoding probability $p$ is improved by considering shared randomness [53], shared entanglement [54], and $d$-level system [55, 49]. The information that can be decoded is not limited to the $m$ bits that are encoded, but also extends to functions that use $k$-out-of-$m$ bits [56, 57]. Using this extension, QRACs have also been applied in machine learning to efficiently encode discrete features [58].
6 Conclusion
In this study, we proposed a recursive and quantum-relaxed algorithm to solve the MAX-CUT problem. The quantum relaxation is achieved using the QRAC, which is incorporated into the existing recursive MAX-CUT algorithm. To show the validity of the proposed algorithm, new properties of the quantum-relaxed Hamiltonian were revealed. According to these findings, we showed that the MAX-CUT problem can be approximately reduced to the problem finding the variable with the highest probability, which is determined by the quantum state optimized through the quantum-relaxed Hamiltonian. We confirmed that the proposed algorithm is consistent with this objective and includes the existing methods as special cases. The numerical experiments on graphs with several hundred nodes showed that the proposed algorithm outperforms the GW algorithm, which is the best-known classical approximation algorithm, on graphs with more than 300 nodes both in the approximation ratio and the empirical time complexity.
Acknowledgement
This work is supported by MEXT Quantum Leap Flagship Program Grant Number JPMXS0118067285 and JPMXS0120319794, and JSPS KAKENHI Grant Number 20H05966. We would like to thank Yohichi Suzuki and Michihiko Sugawara of Keio Quantum Computing Center and Hiroshi C. Watanabe of Kyushu University for discussing the theoretical and experimental justification of RQRAO. RR would like to acknowledge Prof. Hiroshi Imai for information on Gset instances and MAX-CUT .
Appendix A Classical Shadow with Random Magic Measurements
Let us define $(m,1)$-random magic measurements, which is referred to as magic state rounding in [11]. In this section, only $(3,1)$-random magic measurements are introduced, as $(2,1)$- and $(1,1)$-random magic measurements are easily derived similarly.
Definition 3 (Random magic measurements [11])
Let $\mu^{[q]\pm}_{i}$ be the single qubit $(3,1)$-magic state corresponding to the $q$th qubit defined as follows:
${\mu_{1}^{[q]+}:=\mu_{3}^{[q]}(0,0,0),\ \ \ \mu_{1}^{[q]-}:=\mu_{3}^{[q]}(1,1,1),}$
${\mu_{2}^{[q]+}:=\mu_{3}^{[q]}(0,1,1),\ \ \ \mu_{2}^{[q]-}:=\mu_{3}^{[q]}(1,0,0),}$
${\mu_{3}^{[q]+}:=\mu_{3}^{[q]}(1,0,1),\ \ \ \mu_{3}^{[q]-}:=\mu_{3}^{[q]}(0,1,0),}$
${\mu_{4}^{[q]+}:=\mu_{3}^{[q]}(1,1,0),\ \ \ \mu_{4}^{[q]-}:=\mu_{3}^{[q]}(0,0,1),}$
where $\mu_{3}^{[q]}(b^{[q]}_{X},b^{[q]}_{Y},b^{[q]}_{Z})$ is defined as Eq. (11). Random magic measurement is the protocol that obtains ${\bf b}\in\{0,1\}^{3n}$ as follows:
  1. 1.
    Select $i\in\{1,2,3,4\}$ uniformly at random for each $q$.
  2. 2.
    Measure the state $\rho$ in the basis $\{\mu^{[q]\pm}_{i}\}$ for all qubits $q\in[n]$ with the selected $i$.
  3. 3.
    Convert the measurement outcomes to the corresponding classical bits as listed in Table 3.
Table 3: Measurement outcome of the random magic measurement and corresponding classical bits. Here, $i_{q}$ is the selected $i\in\{1,2,3,4\}$ for qubit $q$.
$i_{q}$ Outcome $\big{(}b^{[q]}_{X},b^{[q]}_{Y},b^{[q]}_{Z}\big{)}$
1 0 $(0,0,0)$
1 1 $(1,1,1)$
2 0 $(0,1,1)$
2 1 $(1,0,0)$
3 0 $(1,0,1)$
3 1 $(0,1,0)$
4 0 $(1,1,0)$
4 1 $(0,0,1)$
The bit-strings ${\bf b}$ obtained by a random magic measurement follows the probability distribution $\mathcal{P}_{\rm f}({\bf b})$ (see the proof of Theorem 2).
Because the expectation of the $k$-local Pauli observable in Eq. (20) is written as the expectation of the random variable
$\hat{o}:=3^{k/2}(-1)^{\sum^{k}_{i=1}b^{[q^{(i)}]}_{P}},$ (S.1)
the classical shadow [59] based on random magic measurements can be considered. That is, the expectation value of the $k$-local Pauli observables can be estimated by averaging $\hat{o}$ computed from the corrected bits $\{b^{[q]}_{P}\}$ by the random magic measurements. The upper bound of the variance of random variable $\hat{o}$ is readily derived as
${\rm Var}[\hat{o}]=\mathbb{E}[\hat{o}^{2}]-(\mathbb{E}[\hat{o}])^{2}\leq\mathbb{E}[\hat{o}^{2}]=3^{k},$ (S.2)
which is equivalent to that of random Pauli measurements [59]. The classical snapshot based on random magic measurement is
$\hat{\rho}=\bigotimes^{n-1}_{q=0}\Big{(}3U^{[q]\dagger}|\hat{b}^{[q]}\rangle\langle\hat{b}^{[q]}|U^{[q]}-I\Big{)},\ \ \ U^{[q]}\sim\mathcal{U}_{\rm ms},$ (S.3)
where $\hat{b}^{[q]}$ is the measurement outcome of the $q$th qubit. Equation (S.3) is the same as that of random Pauli measurements except for the unitary ensemble $\mathcal{U}_{\rm ms}$. In the case of random magic measurements, the unitary ensemble is $\mathcal{U}_{\rm ms}=\{U_{i}\}^{4}_{i=1}$, where $\mu_{i}^{+}:=U_{i}|0\rangle\langle 0|U_{i}^{\dagger}$ and $\mu_{i}^{-}:=U_{i}|1\rangle\langle 1|U_{i}^{\dagger}$. Note that $U_{1}$ can be expressed as $U_{1}^{\dagger}=e^{{\rm i}tX}e^{{\rm i}sZ}$ where $s=\pi/8$ and $t={\rm arccos}(\sqrt{(1+1/\sqrt{3})/2})$, and $U_{2}^{\dagger}=U^{\dagger}_{1}X$, $U^{\dagger}_{3}=U^{\dagger}_{1}Y$, $U^{\dagger}_{4}=U^{\dagger}_{1}Z$ [11].
Figure S.1: Accuracy comparison of a classical shadow obtained using random Pauli measurements and that obtained using $(3,1)$-random magic measurements. (a) Absolute error of the expectation value of a $2$-local Pauli observable in a ten qubit system; (b) Fidelity between the true state and the one obtained using state tomography in a five qubit system.
To demonstrate the performance of random magic measurements, two numerical experiments were carried out.
Expectation Estimation
A state $\rho$ is sampled from ten-qubit Haar’s random states and is used throughout the experiments. A hundred 2-local Pauli observables are randomly generated and their expectations are evaluated using random Pauli measurements and random magic measurements, respectively. Although the expectation of the $k$-local Pauli observable is within $[-1,1]$, the range of the expectation estimated by random magic measurements is $[-3^{k/2},3^{k/2}]$. To mitigate the anomaly estimation, the expectation value estimated by the random magic measurements is truncated so that the value is within $[-1,1]$. In Fig. S.1, the accuracy of the expectation evaluation of 2-local Pauli observable is shown. The dependence of the accuracy of the expectation evaluation on the number of shots is almost the same for both the random Pauli measurements and random magic measurements, which is as expected because the variance upper bounds of these methods are the same.
State Tomography
A state $\rho$ is sampled from five-qubit Haar’s random states and used throughout the experiments. $N$ snapshots, $\{\hat{\rho}_{i}\}^{N}_{i=1}$, are collected using random Pauli measurements and the random magic measurements. The candidate matrix is then defined as $\hat{\rho}=\sum^{N}_{i=1}\hat{\rho}_{i}/N$. Here, in general, $\hat{\rho}$ does not satisfy the property of the density matrix, i.e., ${\rm tr}(\hat{\rho})=1$ and $\hat{\rho}\succeq 0$. To avoid improper estimation of the fidelity, the nearest state to $\hat{\rho}$ that satisfied the property of the density matrix is obtained by the optimization, and is defined as the result of the state tomography. The optimization process is as follows. The $n$-qubit nearest state is parametrized as $|\hat{\psi}({\boldsymbol{\theta}})\rangle_{k}=f_{k}/\sqrt{\sum^{2^{n}}_{k^{\prime}=1}f^{*}_{k^{\prime}}f_{k^{\prime}}}$ where $f_{k}=\theta^{\rm Re}_{k}+{\rm i}\theta^{\rm Im}_{k}$ and ${\boldsymbol{\theta}}=\{\theta^{\rm Re}_{k},\theta^{\rm Im}_{k}\}^{2^{n}}_{k=1}$. Here, $|\hat{\psi}({\boldsymbol{\theta}})\rangle_{k}$ denotes the $k$th component of $|\hat{\psi}({\boldsymbol{\theta}})\rangle$ in the computational basis. The objective function is defined as $\mathcal{L}({\boldsymbol{\theta}}):=\|\hat{\rho}-|\hat{\psi}({\boldsymbol{\theta}})\rangle\langle\hat{\psi}({\boldsymbol{\theta}})|\|_{\rm F}$ where $\|\bullet\|_{\rm F}$ is the Frobenius norm. Parameter ${\boldsymbol{\theta}}$ is optimized by L-BFGS to minimize $\mathcal{L}({\boldsymbol{\theta}})$. Then, the tomography result is defined as $\overline{\rho}:=|\hat{\psi}({\boldsymbol{\theta}}^{*})\rangle\langle\hat{\psi}({\boldsymbol{\theta}}^{*})|$, where ${\boldsymbol{\theta}}^{*}$ is the optimized parameter. According to Fig. S.1, the dependence of the state tomography accuracy on the number of shots is almost the same for both the random Pauli measurements and random magic measurements.
Appendix B Node Isolation Problem
Figure S.2: Schematic illustration of the node isolation process. (a) The red and green nodes represent the node to be deleted and the node to be kept, respectively. The red edge indicates the candidate edge. The ensemble edge energy of the candidate edge $\mathcal{E}$ is negative. (b) After removing the candidate edge. (c) Multiplying the sign of the ensemble edge energy of the candidate edge and reconnecting it to the green node. (d) After removing the zero-weight edge.
When the edge weights of problem graph $G$ are integers, the number of isolated nodes, which are defined as nodes with undetermined parity with respect to other nodes, will increase as the iteration progresses. An example is illustrated in Fig. S.2. In Fig. S.2, the node to be deleted and the node to be kept are shown in red and green, respectively. The red edge indicates the candidate whose ensemble edge energy $\mathcal{E}$ is negative. After removing the candidate edge (Fig. S.2), the edge that was connected to the red node at one end is reconnected to the green node (Fig. S.2). At this time, the edge weight is multiplied by the sign of $\mathcal{E}$. As a result, the weight of the edge the bottom of the illustration is updated to zero (Fig. S.2), which means that the black node becomes isolated. If there are $N_{\rm iso}$ isolated nodes after the parity determination procedure, a brute-force search must be carried out on the $M+N_{\rm iso}$ nodes, which is intractable when there are several tens of isolated nodes. One easy way to mitigate the node isolation problem is to apply a small perturbation to all edge weights at random. When perturbed edge weights are used, the probability that a reconnected edge has zero weight becomes negligibly small.
Appendix C Proofs
  • Proof of Theorem 2.
    Consider a quantum channel $\mathcal{E}$ that measures an $n$ qubit state $\rho$ with the measurement basis $\mu_{m}({\bf b})$, which is an $n$ qubit product state of magic states selected uniformly at random:
    $\mathcal{E}(\rho):=\frac{1}{2^{(m-1)n}}\sum_{{\bf b}\in\{0,1\}^{mn}}{\rm tr}(\mu_{m}({\bf b})\rho)\mu_{m}({\bf b}),$ (S.4)
    where $2^{(m-1)n}$ is the total number of $(m,1)$-magic state bases living in $n$-qubit quantum state space. According to [11], the following holds:
    ${\rm tr}\big{(}P^{\otimes k}\mathcal{E}(\rho)\big{)}=\frac{1}{m^{k}}{\rm tr}\big{(}P^{\otimes k}\rho\big{)}$ (S.5)
    In addition, it is easily confirmed that
    ${\rm tr}(P^{\otimes k}\mu_{m}({\bf b}))=\frac{(-1)^{\sum^{k}_{i=1}b^{[q^{(i)}]}_{P}}}{m^{k/2}}.$ (S.6)
    Substituting Eqs. (S.4) and (S.6) into Eq. (S.5) yields
    ${{\rm tr}(P^{\otimes k}\rho)=m^{k}{\rm tr}(P^{\otimes k}\mathcal{E}(\rho))}$
    ${=m^{k}{\rm tr}\Bigg{(}P^{\otimes k}\frac{1}{2^{(m-1)n}}\sum_{{\bf b}\in\{0,1\}^{mn}}{\rm tr}(\mu_{m}({\bf b})\rho)\mu_{m}({\bf b})\Bigg{)}}$
    ${=\frac{m^{k}}{2^{(m-1)n}}\sum_{{\bf b}\in\{0,1\}^{mn}}{\rm tr}(P^{\otimes k}\mu_{m}({\bf b}))f_{m}({\bf b},\rho)}$
    $=\frac{m^{k/2}}{2^{(m-1)n}}\sum_{{\bf b}\in\{0,1\}^{mn}}(-1)^{\sum^{k}_{i=1}b^{[q^{(i)}]}_{P}}f_{m}({\bf b},\rho),$ (S.7)
    where
    $f_{m}({\bf b},\rho):={\rm tr}(\mu_{m}({\bf b})\rho)$ (S.8)
    is a fidelity of $\mu_{m}({\bf b})$ and $\rho$. Meanwhile, it is easily confirmed that
    $1={\rm tr}(\mathcal{E}(\rho))=\frac{1}{2^{(m-1)n}}\sum_{{\bf b}^{\prime}\in\{0,1\}^{mn}}f({\bf b}^{\prime},\rho).$ (S.9)
    Then, the following holds:
    ${{\rm tr}(P^{\otimes k}\rho)=m^{k/2}\,\frac{\sum_{{\bf b}\in\{0,1\}^{mn}}(-1)^{\sum^{k}_{i=1}b^{[q^{(i)}]}_{P}}f_{m}({\bf b},\rho)}{\sum_{{\bf b}^{\prime}\in\{0,1\}^{mn}}f_{m}({\bf b}^{\prime},\rho)}}$
    $=m^{k/2}\,\mathbb{E}_{{\bf b}\sim\mathcal{P}m({\bf b};\rho)}\Big{[}(-1)^{\sum^{k}_{i=1}b^{[q^{(i)}]}_{P}}\Big{]},$ (S.10)
    where
    $\mathcal{P}_{m}({\bf b};\rho):=\frac{f_{m}({\bf b},\rho)}{\sum_{{\bf b}^{\prime}\in\{0,1\}^{mn}}f_{m}({\bf b}^{\prime},\rho)}$ (S.11)
    $\Box$
  • Proof of Corollary 1.
    Consider that a binary variable of node $j$, $b_{j}$, is mapped to $(q^{(j)},P^{(j)})$, where $q^{(j)}$ is a qubit index, $P^{(j)}\in\mathfrak{P}\subset\{X,Y,Z\}$ and $|\mathfrak{P}|=m\in\{1,2,3\}$ (see Sec. 2.3). Equation (S.7) in the case of $k=2$ reduces
    ${\rm tr}\big{(}P^{(j)}_{q^{(j)}}P^{(k)}_{q^{(k)}}\rho\big{)}=m\,\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}(-1)^{b_{j}+b_{k}}\big{]}.$ (S.12)
    The left-hand side of Eq. (S.12) is written as ${\rm tr}\big{(}P_{\langle j\rangle}P_{\langle k\rangle}\rho\big{)}$ in the main body. Substituting the $(m,1)$-QRAC Hamiltonian (Eq. (10)) into Eq. (S.12) leads to
    ${\rm tr}(H_{m}\,\rho)=\sum_{(j,k)\in E}\frac{{\rm tr}(\rho)-m{\rm tr}(P_{\langle j\rangle}P_{\langle k\rangle}\,\rho)}{2}w_{jk}=\sum_{(j,k)\in E}\frac{1-m^{2}\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}(-1)^{b_{j}+b_{k}}\big{]}}{2}w_{jk}.$ (S.13)
    According to the definition of $\texttt{CW}({\bf b})$ (Eq. (2)),
    $\sum_{(j,k)\in E}\frac{(-1)^{b_{j}+b_{k}}}{2}w_{jk}=\sum_{(j,k)\in E}\frac{1}{2}w_{jk}-\texttt{CW}({\bf b}).$ (S.14)
    Substituting Eq. (S.14) into (S.13) gives
    ${\rm tr}(H_{\rm QRAC}\,\rho)=m^{2}\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})]-\frac{m^{2}-1}{2}\sum_{(j,k)\in E}w_{jk}.$ (S.15)
    $\Box$
  • Proof of Corollary 2.
    According to Eq. (10),
    $\begin{array}[]{ll}\displaystyle{H^{2}_{m}}&\displaystyle{=\Big{(}\sum_{(j,k)\in E}\frac{I-mP_{\langle j\rangle}P_{\langle k\rangle}}{2}w_{jk}\Big{)}\cdot\Big{(}\sum_{(j^{\prime},k^{\prime})\in E}\frac{I-mP_{\langle j^{\prime}\rangle}P_{\langle k^{\prime}\rangle}}{2}w_{j^{\prime}k^{\prime}}\Big{)}}\\[14.22636pt] &\displaystyle{=\frac{m^{2}}{4}\sum_{(j,k)\in E}\sum_{(j^{\prime},k^{\prime})\in E}P_{\langle j\rangle}P_{\langle k\rangle}P_{\langle j^{\prime}\rangle}P_{\langle k^{\prime}\rangle}w_{jk}w_{j^{\prime}k^{\prime}}-\frac{mW}{2}\sum_{(j,k)\in E}P_{\langle j\rangle}P_{\langle k\rangle}w_{jk}}+\frac{1}{4}W^{2}\end{array},$
    where $W:=\sum_{(j,k)\in E}w_{jk}$. The expectation value of $H^{2}_{m}$ for an arbitrary state $\rho$ is calculated as
    $\begin{array}[]{ll}\displaystyle{{\rm tr}(H^{2}_{m}\rho)}&\displaystyle{=\frac{m^{2}}{4}\sum_{(j,k)\in E}\sum_{(j^{\prime},k^{\prime})\in E}{\rm tr}(P_{\langle j\rangle}P_{\langle k\rangle}P_{\langle j^{\prime}\rangle}P_{\langle k^{\prime}\rangle}w_{jk}w_{j^{\prime}k^{\prime}}\rho)-\frac{mW}{2}\sum_{(j,k)\in E}{\rm tr}(P_{\langle j\rangle}P_{\langle k\rangle}w_{jk}}\rho)+\frac{1}{4}W^{2}\\[14.22636pt] &\displaystyle{=\frac{m^{2}}{4}\sum_{(j,k)\in E}\sum_{(j^{\prime},k^{\prime})\in E}{\rm tr}(P_{\langle j\rangle}P_{\langle k\rangle}P_{\langle j^{\prime}\rangle}P_{\langle k^{\prime}\rangle}w_{jk}w_{j^{\prime}k^{\prime}}\rho)+W{\rm tr}(H_{m}\rho)-\frac{1}{4}W^{2}}.\end{array}$
    Meanwhile, from Eq. (22),
    $\begin{array}[]{ll}\displaystyle{(\texttt{CW}({\bf b}))^{2}}&\displaystyle{=\Big{(}\sum_{(j,k)\in E}\frac{1-m^{2}(-1)^{b_{j}+b_{k}}}{2m^{2}}w_{jk}+\frac{m^{2}-1}{2m^{2}}W\Big{)}\Big{(}\sum_{(j^{\prime},k^{\prime})\in E}\frac{1-m^{2}(-1)^{b_{j^{\prime}}+b_{k^{\prime}}}}{2m^{2}}w_{j^{\prime}k^{\prime}}+\frac{m^{2}-1}{2m^{2}}W\Big{)}}\\[14.22636pt] &\displaystyle{=\frac{1}{4}\sum_{(j,k)\in E}\sum_{(j^{\prime},k^{\prime})\in E}(-1)^{b_{j}+b_{k}+b_{j^{\prime}}+b_{k^{\prime}}}w_{jk}w_{j^{\prime}k^{\prime}}-\frac{1}{2}W\sum_{(j,k)\in E}(-1)^{b_{j}+b_{k}}w_{jk}+\frac{1}{4}W^{2}}.\end{array}$
    Using Theorem 2 and substituting the expression of ${\rm tr}(H^{2}_{m}\rho)$ derived above yields
    $\begin{array}[]{ll}\displaystyle{\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}(\texttt{CW}({\bf b}))^{2}\big{]}}&\displaystyle{=\frac{1}{m^{4}}{\rm tr}(H^{2}_{m}\rho)+\frac{m^{2}-1}{m^{4}}W{\rm tr}(W\rho)+\frac{(m^{2}-1)^{2}}{4m^{2}}W^{2}}.\end{array}$
    Combining with Eq. (22), the variance of the cut weight can be written as
    $\begin{array}[]{ll}\displaystyle{{\rm Var}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}}&\displaystyle{=\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})^{2}\big{]}-\mathbb{E}_{{\bf b}\sim\mathcal{P}_{m}({\bf b};\rho)}\big{[}\texttt{CW}({\bf b})\big{]}^{2}}\\[14.22636pt] &\displaystyle{=\frac{1}{m^{4}}{\rm tr}(H_{m}^{2}\rho)+\frac{m^{2}-1}{m^{4}}W{\rm tr}(H_{m}\rho)+\frac{(m^{2}-1)^{2}}{4m^{4}}W^{2}}\\[5.69054pt] &\displaystyle{{\hskip 28.45274pt}-\Big{(}\frac{1}{m^{2}}{\rm tr}(H_{m}\rho)+\frac{m^{2}-1}{2m^{2}}W\Big{)}^{2}}\\[14.22636pt] &\displaystyle{=\frac{1}{m^{4}}\Big{(}{\rm tr}(H_{m}^{2}\rho)-{\rm tr}(H_{m}\rho)^{2}\Big{)}}.\end{array}$
    $\Box$
  • Proof of Corollary 3.
    According to Eq. (20),
    $\begin{array}[]{ll}\displaystyle{\mathcal{E}_{j}}&\displaystyle{:={\rm tr}(P_{\langle j\rangle}\rho)}\\[5.69054pt] &\displaystyle{=\sqrt{m}\,\mathbb{E}_{{\bf b}\sim\mathcal{P}_{\rm f}({\bf b};\rho)}[(-1)^{b_{j}}]}\\[5.69054pt] &\displaystyle{=\sqrt{m}\,\big{[}\mathcal{P}(b_{j}=0)\cdot(+1)+\mathcal{P}(b_{j}=1)\cdot(-1)\big{]}}\\[5.69054pt] &\displaystyle{=\sqrt{m}\,\big{[}\mathcal{P}(b_{j}=0)-(1-\mathcal{P}(b_{j}=0))\big{]}}\\[5.69054pt] &\displaystyle{=2\sqrt{m}\mathcal{P}(b_{j}=0)-\sqrt{m}}.\end{array}$
    Then,
    ${\mathcal{P}(b_{j}=0)=\frac{1}{2}+\frac{1}{2\sqrt{m}}\mathcal{E}_{j}}$
    ${\mathcal{P}(b_{j}=1)=1-\mathcal{P}(b_{j}=0)=\frac{1}{2}-\frac{1}{2\sqrt{m}}\mathcal{E}_{j}.}$
    Similarly, when $q^{(j)}\neq q^{(k)}$,
    $\begin{array}[]{ll}\displaystyle{\mathcal{E}_{jk}}&\displaystyle{:={\rm tr}(P_{\langle j\rangle}P_{\langle k\rangle}\rho)}\\[5.69054pt] &\displaystyle{=m\,\mathbb{E}_{{\bf b}\sim\mathcal{P}_{\rm f}({\bf b};\rho)}[(-1)^{b_{j}+b_{k}}]}\\[5.69054pt] &\displaystyle{=m\,\big{[}\mathcal{P}(b_{j}=b_{k})\cdot(+1)+\mathcal{P}(b_{j}\neq b_{k})\cdot(-1)\big{]}}\\[5.69054pt] &\displaystyle{=m\,\big{[}\mathcal{P}(b_{j}=b_{k})-(1-\mathcal{P}(b_{j}=b_{k}))\big{]}}\\[5.69054pt] &\displaystyle{=2m\mathcal{P}(b_{j}=b_{k})-m}.\end{array}$
    Then,
    ${\mathcal{P}(b_{j}=b_{k})=\frac{1}{2}+\frac{1}{2m}\mathcal{E}_{jk}}$
    ${\mathcal{P}(b_{j}\neq b_{k})=1-\mathcal{P}(b_{j}=b_{k})=\frac{1}{2}-\frac{1}{2m}\mathcal{E}_{jk}.}$
    However, when $q^{(j)}=q^{(k)}$
    $\begin{array}[]{ll}\displaystyle{\mathcal{P}(b_{j}=b_{k})}&\displaystyle{=\mathcal{P}(b_{j}=0)\cdot\mathcal{P}(b_{k}=0)+\mathcal{P}(b_{j}=1)\cdot\mathcal{P}(b_{k}=1)}\\[5.69054pt] &\displaystyle{=\frac{1}{2}+\frac{1}{2m}\mathcal{E}_{j}\mathcal{E}_{k}}\end{array}$
    and
    $\begin{array}[]{ll}\displaystyle{\mathcal{P}(b_{j}\neq b_{k})}&\displaystyle{=1-\mathcal{P}(b_{j}=b_{k})}\\[5.69054pt] &\displaystyle{=\frac{1}{2}-\frac{1}{2m}\mathcal{E}_{j}\mathcal{E}_{k}}.\end{array}$
    $\Box$
Appendix D Matrix Product State
Let $\Psi\in\mathbb{C}^{2^{n}}$ be $n$-qubit quantum state. Generally, $\Psi$ can be represented as the matrix multiplication of $n$ matrices as
$\Psi=A^{[n-1]}A^{[n-2]}\cdots A^{[1]}A^{[0]},$ (S.16)
where
$A^{[i]}\in\mathbb{C}^{\min(2^{i+1},2^{n-i-1},\chi)\times 2\times\min(2^{i},2^{n-i-1},\chi)}$ (S.17)
and
$(A^{[i]}A^{[j]})_{k(ab)m}=\sum^{\min(2^{i},2^{n-i-1},\chi)}_{l=1}(A^{[i]})_{kal}(A^{[j]})_{lbm}.$ (S.18)
The subscript in the parentheses, $(ab)$, summarizes two indices in one index, i.e., $(a,b)=(1,1),(1,2),\ldots,(2,1),(2,2),\ldots$. The expression of Eq. (S.16) is called the MPS [60]. The dimension between matrices is called the bond dimension and can be restricted to $\chi$ at the expense of representation power. The number of components of an $n$-qubit state is $2^{n}$, whereas the total number of matrix components of MPS is $\sim 2n\chi^{2}$. For example, a $4$-qubit state $|\psi\rangle\in\mathbb{C}^{2^{4}}$ with bond dimension $\chi=2$ can be represented as
${\psi_{(abcd)}=\sum_{\alpha,\beta,\gamma,\delta,\epsilon}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{a}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{a}\end{bmatrix}_{\alpha\beta}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}\end{bmatrix}_{\beta\gamma}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}\end{bmatrix}_{\gamma\delta}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{d}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{d}\end{bmatrix}_{\delta\epsilon}}$
$\approx\sum_{\alpha,\beta,\gamma,\delta,\epsilon}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{a}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{a}\end{bmatrix}_{\alpha\beta}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{b}\end{bmatrix}_{\beta\gamma}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}&\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{c}\end{bmatrix}_{\gamma\delta}\begin{bmatrix}\begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{d}\\[14.22636pt] \begin{bmatrix}\bullet\\ \bullet\end{bmatrix}_{d}\end{bmatrix}_{\delta\epsilon},$ (S.19)
where $\bullet$ indicates the matrix element and the subscripts represent the index symbols of the inner submatrix elements.
Let $P$ be the identity or Pauli matrix and
$H=\sum^{K}_{k=1}C_{k}\bigotimes^{n-1}_{l=0}P_{kl}$ (S.20)
be a $n$-qubit Hamiltonian that consists of a weighted sum of $K$ Pauli products. Similar to the MPS, Eq. (S.20) can be represented as matrix multiplication of $n$ matrices as
$H=B^{[n-1]}B^{[n-2]}\cdots B^{[1]}B^{[0]},$ (S.21)
where
$B^{[i]}\in\left\{\begin{array}[]{ll}\displaystyle{\mathbb{C}^{K\times 2\times 2\times 1}}&i=0\\[14.22636pt] \displaystyle{\mathbb{C}^{1\times 2\times 2\times K}}&i=n-1\\[14.22636pt] \displaystyle{\mathbb{C}^{K\times 2\times 2\times K}}&{\rm else}\end{array}\right.,$ (S.22)
$\left\{\begin{array}[]{ll}\displaystyle{(B^{[i]})_{kaa^{\prime}1}=C_{k}(P_{ki})_{aa^{\prime}}}&i=0\\[14.22636pt] \displaystyle{(B^{[i]})_{1aa^{\prime}k}=(P_{ki})_{aa^{\prime}}}&i=n-1\\[14.22636pt] \displaystyle{(B^{[i]})_{kaa^{\prime}k}=(P_{ki})_{aa^{\prime}}}&{\rm else}\end{array}\right.$ (S.23)
and
$(B^{[i]}B^{[j]})_{l(ab)(a^{\prime}b^{\prime})m}=\sum^{K}_{k=1}(B^{[i]})_{laa^{\prime}k}(B^{[j]})_{kbb^{\prime}m}.$ (S.24)
For example,
${H_{(abcd)(a^{\prime}b^{\prime}c^{\prime}d^{\prime})}=(C_{1}IXYI+C_{2}IZII+C_{3}IYIX)_{(abcd)(a^{\prime}b^{\prime}c^{\prime}d^{\prime})}}$
$=\sum_{\alpha,\beta,\gamma,\delta,\epsilon}\begin{bmatrix}[I]_{aa^{\prime}}&[I]_{aa^{\prime}}&[I]_{aa^{\prime}}\end{bmatrix}_{\alpha\beta}\begin{bmatrix}[X]_{bb^{\prime}}\\ &[Z]_{bb^{\prime}}\\ &&[Y]_{bb^{\prime}}\end{bmatrix}_{\beta\gamma}\begin{bmatrix}[Y]_{cc^{\prime}}\\ &[I]_{cc^{\prime}}\\ &&[I]_{cc^{\prime}}\end{bmatrix}_{\gamma\delta}\begin{bmatrix}[C_{1}I]_{dd^{\prime}}\\ [C_{2}I]_{dd^{\prime}}\\ [C_{3}X]_{dd^{\prime}}\end{bmatrix}_{\delta\epsilon}$ (S.25)
The expression of Eq. (S.21) is called the matrix product operator (MPO) [60].
The MPO is a set of sparse matrices and thus can be efficiently stored in classical memory in COOrdinate (COO), Compressed Sparse Row (CSR), or Compressed Sparse Column (CSC) formats.
Using the MPS and MPO, the expectation value $\langle\Psi|H|\Psi\rangle$ can be efficiently calculated by the matrix multiplication of $\{A^{[i]}\}$ and $\{B^{[i]}\}$.
All edge energy can be simultaneously evaluated by modifying the MPO of the problem Hamiltonian.
$\left\{\begin{array}[]{ll}\displaystyle{(B^{[i]})_{kaa^{\prime}1}=(P_{ki})_{aa^{\prime}}}&i=0\\[14.22636pt] \displaystyle{(B^{[i]})_{kaa^{\prime}k}=(P_{ki})_{aa^{\prime}}}&{\rm else}\end{array}\right.$ (S.26)
For example,
${\begin{bmatrix}C_{1}IXYI_{(abcd)(a^{\prime}b^{\prime}c^{\prime}d^{\prime})}\\ C_{2}IZII_{(abcd)(a^{\prime}b^{\prime}c^{\prime}d^{\prime})}\\ C_{3}IYIX_{(abcd)(a^{\prime}b^{\prime}c^{\prime}d^{\prime})}\end{bmatrix}}$
$=\sum_{\alpha,\beta,\gamma,\delta,\epsilon}\begin{bmatrix}[I]_{aa^{\prime}}\\ &[I]_{aa^{\prime}}\\ &&[I]_{aa^{\prime}}\end{bmatrix}_{\alpha\beta}\begin{bmatrix}[X]_{bb^{\prime}}\\ &[Z]_{bb^{\prime}}\\ &&[Y]_{bb^{\prime}}\end{bmatrix}_{\beta\gamma}\begin{bmatrix}[Y]_{cc}\\ &[I]_{cc^{\prime}}\\ &&[I]_{cc^{\prime}}\end{bmatrix}_{\gamma\delta}\begin{bmatrix}[C_{1}I]_{dd^{\prime}}\\ [C_{2}I]_{dd^{\prime}}\\ [C_{3}X]_{dd^{\prime}}.\end{bmatrix}_{\delta\epsilon}$ (S.27)
Appendix E Implementation Details for GW, CirCut, the RQAOA, and QRAO
For GW, an official MATLAB implementation of SDPT3 [61, 62] is used to solve the SDP. The gap tolerance is set to $10^{-9}$ for high-precision calculation. For hyperplane cutting, which is a post-processing step in GW performed after computing the SDP solution, in-house Python code is used. The number of hyperplane cuttings is set to $10,000$.
For CirCut [16], the official Fortran implementation (https://www.cmor-faculty.rice.edu/~zhang/circut/index.html) is used. The hyperparameters for CirCut are not changed from the default ones.
For RQAOA, in-house Fortran code is used. The analytic form of $\langle\psi_{\rm QAOA}({\boldsymbol{\theta}})|Z_{j}Z_{k}|\psi_{\rm QAOA}({\boldsymbol{\theta}})\rangle$ [3] is used to evaluate the expectation value of the problem Hamiltonian and edge energy, where $|\psi_{\rm QAOA}({\boldsymbol{\theta}})\rangle=e^{i\beta B}e^{i\gamma C}|+\rangle^{\otimes n}$ is a level-1 QAOA ansatz, $B=\sum^{n-1}_{i=0}X_{i}$, $C=\sum_{(j,k)\in E}w_{jk}Z_{j}Z_{k}$, and ${\boldsymbol{\theta}}=\{\beta,\gamma\}$. Only the optimal $\gamma$ is determined using a grid search; the optimal $\beta$ is calculated for each $\gamma$ using the same method used in [14]. We derived the concrete representation of the optimal $\beta$ as
$\beta^{*}=-\frac{1}{4}{\rm arctan}\left(\frac{\frac{1}{2}(R_{11}-R_{22}-R_{33}+R_{44})+R_{14}-R_{23}}{-I_{12}-I_{13}+I_{24}+I_{34}}\right)+\frac{\pi}{8},$ (S.28)
where
$\begin{bmatrix}R_{11}&R_{12}+iI_{12}&R_{13}+iI_{13}&R_{14}+iI_{14}\\ R_{12}-iI_{12}&R_{22}&R_{23}+iI_{23}&R_{24}+iI_{24}\\ R_{13}-iI_{13}&R_{23}-iI_{23}&R_{33}&R_{34}+iI_{34}\\ R_{14}-iI_{14}&R_{24}-iI_{24}&R_{34}-iI_{34}&R_{44}\end{bmatrix}:=\sum_{(u,v)\in E}w_{uv}\rho_{uv}$ (S.29)
and
$\rho_{uv}:={\rm tr}_{uv}\Big{(}e^{i\gamma C}|+\rangle^{\otimes n}\langle+|^{\otimes n}e^{-i\gamma C}\Big{)}.$ (S.30)
Here, ${\rm tr}_{uv}(\sigma)$ denotes the reduced state obtained by tracing out, excepting the $u$th and $v$th qubits from $\sigma$. For an algorithm to calculate $\rho_{uv}$, see [14]. The parameter spaces are set to $\beta\in[0,\frac{\pi}{2}]$ and $\gamma\in[0,\pi]$. Once the optimal $\gamma$ is found in $50$ equidistant grids of $[0,\pi]$, the more precise solution is searched for in the fine-grained grids around the course-grained solution with a mesh size is $1/50$. Fine-graining is repeated twice, resulting in a mesh size of $1/250$ of the parameter space.
For QRAO, in-house Python code is used. The MPS ansatz with bond dimension $\chi=2$ is used for $|\psi({\boldsymbol{\theta}})\rangle$ and Pauli rounding is used for decoding. The same optimizer as used for RQRAO, L-BFGS with the strong Wolfe condition, is used to optimize $|\psi({\boldsymbol{\theta}})\rangle$.
Appendix F Additional Sensitivity Analysis Results (ExpHyp)
Figure S.3: Results of the ablation study. Only one hyperparameter is varied for each trial. The problem graph instances are $3$-regular graphs with ±$\pm 1$ edge weights. For each number of nodes, $10$ graph instances are randomly generated. The varying hyperparameters are the (a) number of ensembles $N$; (b) scale factor $S$; and (c) bond dimension $\chi$.
Here, the sensitivities of all RQRAO hyperparameters except for the number of embeddings per qubit $m$ on the performance of the cut weight and runtime are reported. The results are summarized in Fig. S.3. In brief, a higher cut weight is obtained through a longer runtime.
Except for the number of embeddings per qubit $m$, the sensitive hyperparameters with respect to cut weight performance are the number of ensembles $N$ and scale factor $S$. In Fig. S.3, the cut weight increases as the number of ensembles $N$ increases. In addition, as the number of ensembles increases, the number of edges whose parities are simultaneously determined in one optimization step tends to decrease. This is because the edges, which are a candidate for determining the parity when the number of ensembles is small, drop out as the accuracy of the parity confidence improves, which is achieved by increasing the number of ensembles. This trend (i.e., the number of candidates decreases with as the accuracy of the parity confidence increases), can be seen in the larger scale factor $S$ in Fig. S.3. A larger $S$ tends to decrease the number of candidates because the number of edges with nonzero $\mathcal{E}_{jk}$ decreases.
The effect of the bond dimension $\chi$ on the cut weight performance needs to be carefully considered. Recall that a large entanglement can be treated using a large bond dimension. The $\chi=1$ result is worse both in terms of relative cut weight and runtime than the $\chi>1$ results. In the case of $\chi=1$, $|\psi({\boldsymbol{\theta}})\rangle$ is restricted to a direct product state that tends to be trapped in a lower energy state than in the $\chi>1$ case because the search space is much smaller. As a result, each edge energy in an ensemble, $\mathcal{E}^{(t)}_{jk}$, is no longer biasedly distributed in either positive or negative sides because the lower energy state includes smaller information of a large cut weight. By contrast, when $\chi>1$, no significant difference in the relative cut weight can be seen. This does not mean, however, that a large entanglement will not affect the cut weight performance because even the value of $\chi=10$, which was the maximum value considered in the experiment, is small compared with the maximum bond dimension that the system can take, which is $\chi_{\rm max}=2^{\lfloor n/2\rfloor}$. For example, consider the case of $|V|=300$. The number of qubits is assumed to be $300/3=100$, which can be achieved by a $(3,1)$-QRAC formulation if the graph is sufficiently sparse. In this case, $\chi_{\rm max}=2^{\lfloor 99/2\rfloor}\approx 10^{14}$, which is much larger than the experiments considered in this study. To take into account a large bond dimension within a reasonable runtime, computation using a real quantum device would be required.